### v1

# Learning High-Order Relationships of Brain Regions

Abstract
Discovering reliable and informative interactions among brain regions from functional magnetic resonance imaging (fMRI) signals is essential in neuroscientific predictions of cognition. Most of the current methods fail to accurately characterize those interactions because they only focus on pairwise connections and overlook the high-order relationships of brain regions. We delve into this problem and argue that these high-order relationships should be maximally informative and minimally redundant (MIMR). However, identifying such high-order relationships is challenging and highly under-explored. Methods that can be tailored to our context are also non-existent. In response to this gap, we propose a novel method named HyBRiD that aims to extract MIMR high-order relationships from fMRI data. HyBRiD employs a Constructor to identify hyperedge structures, and a Weighter to compute a weight for each hyperedge. HyBRiD achieves the MIMR objective through an innovative information bottleneck framework named multi-head drop-bottleneck with theoretical guarantees. Our comprehensive experiments demonstrate the effectiveness of our model. Our model outperforms the state-of-the-art predictive model by an average of $12.1\%$, regarding the quality of hyperedges measured by CPM, a standard protocol for studying brain connections.
1 Introduction
Discovering reliable and informative relations among brain regions using fMRI signals is crucial for predicting human cognition in and understanding brain functions (Kucian et al., 2008; 2006; Li et al., 2015b; a; Satterthwaite et al., 2015; Wang et al., 2016).
However, despite the clear multiplexity of the brain’s involvement in cognition (Logue & Gould, 2014; Barrasso-Catanzaro & Eslinger, 2016; Knauff & Wolf, 2010; Reineberg et al., 2022), the current imaging biomarker detection methods (Shen et al., 2017; Gao et al., 2019; Li et al., 2021) focus only on the contributing roles of the pairwise connectivity edges. In contrast, most brain functions involve distributed patterns of interactions among multiple regions (Semedo et al., 2019). For instance, executive planning requires the appropriate communication of signals across many distinct cortical areas (Logue & Gould, 2014). Pairwise connectivity, without considering the brain’s high-order structure, can lead to inconsistent findings across studies with low prediction performance. Although there are few works (Zu et al., 2016; Xiao et al., 2019; Li et al., 2022) working on discovering the high-order relationships of brain regions, they cannot effectively extract meaningful patterns towards a neurological outcome since 1) the discovering process of high-order relationships is not guided by cognitive targets; 2) the weak expressiveness of traditional machine learning methods (e.g. lasso regression) hinders the accurate modeling of brain region interactions (Cao et al., 2022; Richards et al., 2019).
In this paper, we aim to identify high-order relationships that are informative toward a neurological outcome. However, unlike pairwise relations, the number of possible high-order relations is exponential. To identify the most informative ones from the exponential space, we propose our objective: maximally informative and minimally redundant (MIMR). That is to say, we aim to maximize the information contained in the high-order relationships towards a neurological outcome (informativeness) while diminishing the participation of unrelated brain regions (redundant). Such a criterion, on the one hand, ensures the predictive performance of these high-order relationships; on the other hand, it endows the model with the capacity to identify more succinct and interpretable structures (Yu et al., 2020; Miao et al., 2022a; b; Chen & Ying, 2023). A formal definition of the MIMR criterion could be found in Equation 8 from an information bottleneck view.
Figure 1: We identify high-order relationships of brain regions, where hyperedges possess strong relevance to cognition (maximal informativeness). Meanwhile, they contain the least irrelevant information.
We formulate high-order relationships as hyperedges in a hypergraph, where regions are treated as nodes. Unlike a traditional graph where edges connect only two nodes, a hypergraph allows edges, known as hyperedges, to connect any number of nodes. The hypergraph should be weighted, and the weights of hyperedges are considered as strengths of high-order relationships, which contain the information relevant to cognition (Figure 1).
However, current methods for hypergraph construction, which mostly are based on neighbor distances and neighbor reconstruction (Wang et al., 2015; Liu et al., 2017; Jin et al., 2019; Huang et al., 2009), are unsuitable in our context for several reasons: 1) they are unable to learn MIMR hyperedges due to the absence of a tractable objective for learning such hyperedges. 2) they fall short of learning consistent structures across subjects, which contradicts the belief that the cognitive function mechanism of healthy human beings should be similar (Wang et al., 2023). 3) the number of hyperedges is restricted to the number of nodes, which may lead to sub-optimal performance. Furthermore, although information bottleneck (IB) is a prevailing solution to learn MIMR representations in deep learning (Kim et al., 2021; Alemi et al., 2016; Luo et al., 2019), existing IB methods focus on extracting compressed representations of inputs instead of identifying underlying structures such as hypergraphs. Harnessing the IB framework for identifying hypergraphs necessitates both architectural innovations and theoretical derivations.
Proposed work
In this paper, we propose Hypergraph of Brain Regions via multi-head Drop-bottleneck (HyBRiD), a novel approach for identifying maximally informative yet minimally redundant high-order relationships of brain regions. The overall pipeline of HyBRiD is depicted in Figure 2. HyBRiD is equipped with a Constructor and a Weighter. The Constructor identifies the hyperedge structures of brain regions by learning sets of masks, and the Weighter computes a weight for each hyperedge. To advance the IB principle for hyperedge identification, we further propose multi-head drop-bottleneck and derive its optimization objective.
HyBRiD gets rid of searching in an exponential space through learning masks to identify hyperedges, which guarantees computational efficiency. Its feature-agnostic masking mechanism design ensures HyBRiD to learn consistent structures across subjects. Moreover, the model is equipped with a number of parallel heads, and each head is dedicated to a hyperedge. Through this, HyBRiD is able to identify any number of hyperedges, depending on how many heads it is equipped with. Additionally, the proposed multi-head drop-bottleneck theoretically guarantees the maximal informativeness and minimal redundancy of the identified hyperedges.
We evaluate our methods on $8$ datasets of human subjects under different conditions and at different participation times. We quantitatively evaluate our approach by a commonly used protocol for studying brain connections, CPM (Shen et al., 2017) (Appendix B), and show that our model outperforms the state-of-the-art deep learning models by an average of $12.1\%$ on a comprehensive benchmark. Our post-hoc analysis demonstrates that hyperedges of higher degrees are considered more significant, which indicates the significance of high-order relationships in human brains.
2 Problem Definition & Notations
Input
We consider ABCD, which is a widely used benchmark in fMRI studies (5.1). This dataset is a collection of instances, where each instance corresponds to an individual human subject. An instance is represented by the pair $(X,Y)$. $X\in\mathbb{R}^{N\times d}$ represents the features for each subject, where $N$ is the number of brain regions and $d$ is the number of feature dimensions. The features are obtained by connectivity, consistent with previous works (5.1). $Y\in\mathbb{R}$ denotes the prediction target, such as fluid intelligence. Section 5 elaborates more details about datasets and preprocessing procedures. For how to obtain features of regions from raw data, see Paragraph 5.1 for more details.
Goal
Based on the input $X$, HyBRiD aims to learn a weighted hypergraph of the brain, where regions are nodes. To achieve this, HyBRiD identifies a collection of hyperedges $H=(\bm{h}^{1},\bm{h}^{2},\cdots,\bm{h}^{K})$, and assigns weights $\bm{w}=[w^{1},w^{2},\cdots,w^{K}]^{T}$ for all hyperedges. These hyperedges and their weights, which represent strengths of hyperedges, are expected to be the most informative (i.e. relevant to the target $Y$) yet the least redundant.
Representation of hyperedges
As mentioned before, we use $H$ to denote the collection of hyperedge structures and $\bm{h}^{k}$ to denote the $k$-th hyperedge. To associate hyperedges with node memberships, we use the following representation for a hyperedge:
$\bm{h}^{k}=\bm{m}^{k}\odot X\in\mathbb{R}^{N\times d}$ (1)
where $\bm{m}^{k}\in\{0,1\}^{N}$ is a mask vector and $\odot$ denotes broadcasting element-wise multiplication. In other words, each $\bm{h}^{k}$ is a random row-zeroed version of $X$.
3 Related Work
Hypergraph construction
Existing hypergraph construction methods are mostly based on neighbor reconstruction and neighbor distances. For example, the $k$ nearest neighbor-based method (Huang et al., 2009) connects a centroid node and its $k$ nearest neighbors in the feature space to form a hyperedge. Wang et al. (2015); Liu et al. (2017); Jin et al. (2019); Xiao et al. (2019) further refine these neighbor connections through various regularization. However, the number of hyperedges of these methods is restricted to the number of nodes, and hyperedges obtained by these methods are inconsistent across instances. Zhang et al. (2022; 2018) proposed to iteratively refine a noisy hypergraph, which is obtained by the aforementioned methods. Therefore, they share the same limitations as the aforementioned methods. In addition, these methods are unable to learn MIMR hyperedges due to the absence of a tractable objective. Other methods such as attributed-based methods (Huang et al., 2015; Joslyn et al., 2019) are ill-adapted to our context since they require discrete labels or a prior graph topology. Different from these methods, we provide a way to learn a consistent hypergraph through a deep-learning model without any prior topology. Furthermore, thanks to the proposed multi-head drop-bottleneck, these hyperedges are theoretically ensured MIMR.
High-order relationships in fMRI
Although some methods are working on high-order relationships in fMRI, they are limited or inconsistent with our MIMR objective. Xiao et al. (2019); Li et al. (2022) used the existing non-learning-based hypergraph construction methods, which may lead to noisy and inexpressive hypergraphs. Zu et al. (2016); Santoro et al. (2023) enumerated all hyperedges with degrees lower than 3, which can only discover a tiny portion of all possible hyperedges in exponential space and is not scalable to a large degree. Rosas et al. (2019) proposed O-information, which reflects the balance between redundancy and synergy. The O-information metric is utilized by Varley et al. (2023) to study fMRI data. However, the objective of these methods is not consistent with ours: although both of us are quantifying the redundancy of high-order relations, our method is to learn those that are most informative toward a cognition score, while theirs is to depict the synergy and redundancy within a system.
Information bottleneck
Information bottleneck (IB) (Tishby et al., 2000) is originally a technique in data compression. The key idea is to extract a summary of data, which contains the most relevant information to the objective. Alemi et al. (2016) first employed an IB view of deep learning. After that, IB has been widely used in deep learning. The applications span areas such as computer vision (Luo et al., 2019; Peng et al., 2018), reinforcement learning (Goyal et al., 2019; Igl et al., 2019), natural language processing (Wang et al., 2020) and graph learning (Yu et al., 2020; 2022; Xu et al., 2021; Wu et al., 2020). Unlike these studies that use IB to extract a compressed representation or a select set of features, our approach focuses on identifying the underlying structures of the data.
Connectivity-based cognition prediction
Recently, deep learning techniques have been increasingly employed in predicting cognition based on the connectivity of brain regions. A substantial portion of these studies (Ahmedt-Aristizabal et al., 2021; Li et al., 2019; Cui et al., 2022b; Kan et al., 2022a; Cui et al., 2022a; Said et al., 2023) model the brain network as a graph, in which regions act as nodes and pairwise correlations form the edges. These methods predominantly utilize Graph Neural Networks (GNNs) to capture the connectivity information for predictions. In addition to GNNs, Kan et al. (2022b) proposed to use transformers with a specially designed readout module, leveraging multi-head attention mechanisms to capture pairwise connectivity. However, these methods heavily rely on pairwise connectivity and neglect more intricate higher-order relationships. This oversight, on the one hand, leads to sub-optimal prediction performances; and on the other hand, prevents domain experts from acquiring insightful neuroscience interpretations, given that human cognition often involves multiple regions.
4 Method
Method Overview
HyBRiD consists of a Constructor $\mathcal{F}_{c}$, a Weighter $\mathcal{F}_{w}$, and a LinearHead $\mathcal{F}_{l}$. At a high level, the Constructor $\mathcal{F}_{c}$ is responsible for identifying hyperedges $H$ from the data to construct the hypergraph. After that, the Weighter $\mathcal{F}_{w}$ calculates a weight for each hyperedge. Finally, based on all the weights $\bm{w}$, the LinearHead $\mathcal{F}_{l}$ predicts the label $Y$. An illustration of this pipeline is shown in Figure 2. The pipeline can be formulated as
[…]
We will elaborate on all of the details of the architecture below.
Figure 2: Overview of the HyBRiD pipeline when the total number of hyperedges $K=3$. hyperedge are in distinct colors for clarity. The Constructor identifies hyperedges in the hypergraph, where regions are nodes. The Weighter computes a weight for each hyperedge. These weights, representing strengths of hyperedges, are expected to be informative in terms of our target $Y$. There are two separate phases after obtaining weights of hyperedges: 1) Training. The model’s parameters are trained under the supervision of $Y$; 2) Evaluation. The output weights, as well as pairwise weights, are fed into the CPM (see Appendix B).
4.1 Learning the hypergraph by Multi-head Masking
Given an instance, represented by $X=[X_{1},X_{2},\cdots,X_{N}]^{T}\in\mathbb{R}^{N\times d}$, where $X_{i}\in\mathbb{R^{d}}$ is a column vector representing the features of region $i$. These regions are nodes in the hypergraph we are going to construct. Hyperedges in the hypergraph can be beneficial in the learning task because it is essential to model the interactions between more than two regions.
Hyperedges construction
In this paragraph, we elaborate how the Constructor identifies the hyperedges, i.e. $H=\mathcal{F}_{c}(X)$.
Suppose the number of hyperedges is $K$, which is a predefined hyperparameter. We assign a head to each hyperedge. Each head is responsible for constructing a hyperedge by selecting nodes belonging to that hyperedge.
Specifically, to construct the $k$-th hyperedge, the Constructor’s $k$-th head outputs a column vector $\bm{m}^{k}\in\{0,1\}^{N}$, where each element in the vector corresponds to a region.
$\bm{m}^{k}$ $=[\mathbbm{1}(p^{k}_{\theta,1}),\mathbbm{1}(p^{k}_{\theta,2}),\cdots,\mathbbm{1}(p^{k}_{\theta,N})]^{T}\in\{0,1\}^{N}$ (3)
where $p^{k}_{\theta,i},i=1,2,\cdots,N$ are learnable parameters. $\mathbbm{1}:[0,1]\mapsto\{0,1\}$ is an indicator function, which is defined as $\mathbbm{1}(x)=1$ if $x>0.5$ and $\mathbbm{1}(x)=0$ if $x\leq 0.5$. And $\bm{m}^{k}$ is a column vector corresponding to the $k$-th hyperedge. Note that since there is no gradient defined for the indicator operation, we employ the stopgradient technique (Oord et al., 2017; Bengio et al., 2013) to approximate the gradient.
Figure 3: Architecture details of the Constructor and the Weighter when the number of nodes $N=6$ and the number of hyperedges $K=3$. At a high level, the Constructor learns the hyperedge structure by masking nodes. The Weighter computes the weight of each hyperedge based on the the hyperedge’s member nodes and their features.
In the vector $\bm{m}^{k}$, 00 indicates nodes that are masked out, and $1$ indicates nodes that are not masked. Nodes that are not masked are considered to form a hyperedge together. We use $\bm{h}^{k}$ to represent the masked version of $X$
$\bm{h}^{k}$ $=\bm{m}^{k}\odot X$ (4)
$=[\bm{m}^{k}_{1}X_{1},\bm{m}^{k}_{2}X_{2},\cdots,\bm{m}^{k}_{N}X_{N}]\in\mathbb{R}^{N\times d}$
where $\odot$ is the broadcast element-wise multiplication. $\bm{m}^{k}_{j}$ is the $j$-th element of the vector $\bm{m}^{k}$.
We obtain $K$ hyperedges for $K$ sets of masks. We use $H$ to denote the collection of all hyperedges.
$H=(\bm{h}^{1},\bm{h}^{2},\cdots,\bm{h}^{K})$ (5)
Hyperedge weighting
After obtaining the structure (i.e. member nodes) of each hyperedge, the Weighter will calculate each hyperedge’s weight, which is supposed to indicate the importance of that hyperedge, based on the member nodes and their features, i.e. $\bm{w}=\mathcal{F}_{w}(H)$.
These weights are obtained by a $\operatorname{Readout}$ module, which is composed of: 1) summing over all the non-masked nodes feature-wisely; 2) dim reduction operation.
$w^{k}=\operatorname{Readout}(\bm{h}^{k})=\operatorname{DimReduction}({\bm{m}^{k}}^{T}\bm{h}^{k})\in\mathbb{R}$ (6)
where $w^{k}$ is the weight of the $k$-th hyperedge and $\operatorname{DimReduction}$ is an MLP with ReLU activations, where the output dimension is 1. For all hyperedges, we obtain $K$ hyperedges in total, $\bm{w}=[w^{1},w^{2},\cdots,w^{K}]^{T}\in\mathbb{R}^{K}$. Finally, these weights will be fed into the final linear head to predict the label of the instance,
$\hat{Y}=\mathcal{F}_{l}(\bm{w})\in\mathbb{R}$ (7)
In contrast to previous hypergraph construction methods (Jin et al., 2019; Xiao et al., 2019), which identify hyperedges by refining neighbors and assign weights by aggregating node features, HyBRiD makes these procedures learnable and thus is able to identify MIMR hyperedges in a data-driven way through expressive neural networks. We decide the number of hyperedges $K$ according to the study in Appendix E.2.
4.2 Optimization Framework
We propose a new IB framework named multi-head drop-bottleneck to optimize HyBRiD. To adopt an information bottleneck view of HyBRiD, we consider $X$, $Y$ and $H$ are random variables in the Markovian chain $X\leftrightarrow Y\leftrightarrow H$. According to our MIMR objective, we optimize
$\arg\max I(H;Y)-\beta I(H;X)$ (8)
where $\beta$ is a coefficient trading off informativeness and redundancy. Since optimizing the mutual information for high-dimensional continuous variables is intractable, we instead optimize the lower bound of Equation 8. Specifically, for the first term (informativeness), it is easy to show
$I(H;Y)$ $=\mathbb{H}[Y]-\mathbb{H}[{Y|H}]$ (9)
$=\mathbb{H}[Y]+\mathbb{E}_{p(Y,H)}[\log p(Y|H)]$
$=\mathbb{H}[Y]+\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)]+\mathbb{E}_{p(H)}[\operatorname{KL}(p(Y|H)|q_{\phi}(Y|H))]$
$\geq\mathbb{H}[Y]+\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)]$
where $\mathbb{H}[\cdot]$ is the entropy computation. Since there is no learnable component in the entropy of $Y$, we only need to optimize the second term $\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)]$. $q_{\phi}$ can be considered as a model that predicts $Y$ based on $H$, which essentially corresponds to $\mathcal{F}_{l}\circ\mathcal{F}_{w}$, where $\circ$ is the function composition. In practice, we set $q_{\phi}$ as a Gaussian model with variance $1$ most probabilistic machine learning models for continuous data modeling. (Alemi et al., 2016; Luo et al., 2019; Peng et al., 2018; Kingma & Welling, 2013).
For the second term (redundancy) in Equation 8, we have
Proposition 1.
(Upper bound of $I(H;X)$ in multi-head drop-bottleneck)
$I(H;X)\leq\sum_{k=1}^{K}I(\bm{h}^{k};X)\leq\sum_{k=1}^{K}\sum_{i=1}^{N}I(\bm{h}^{k}_{i};X_{i})=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i})$ (10)
where $\bm{h}^{i}_{j}$ and $X_{j}$ is the $j$-th row of $\bm{h}^{i}$ and $X$ respectively. $p^{k}_{\theta,i}$ is the mask probability in Equation 3. $\mathbb{H}$ is the entropy computation. The equality holds if and only if nodes are independent and hyperedges do not overlap. The second inequality is inspired by Kim et al. (2021). The proof of the proposition can be found in Appendix A.
Therefore, instead of optimizing the intractable objective 8, we optimize its upper bound (i.e. loss function) according to Equation 9 and 10.
$\mathcal{L}$ $=\|Y-\mathcal{F}_{l}\circ\mathcal{F}_{w}\circ\mathcal{F}_{c}(X)\|_{2}^{2}+\beta\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i})\geq-I(H,Y)+\beta I(H,X)$ (11)
The learnable components are the shallow embeddings in Equation 3, the $\operatorname{DimReduction}$ MLP in Equation 6 and the LinearHead $\mathcal{F}_{l}$ in Equation 7. For how do we choose the trade-off coefficient $\beta$, see Appendix E.3 for more discussions.
4.3 Computational Complexity
Suppose we have $N$ regions as $N$ nodes and $K$ hyperedges. For the Constructor, the unignorable computation is from the mask operation. The computational complexity of this step for each hyperedge is $O(Nd)$ since it does a pairwise multiplication operation of a matrix of size $N\times d$. Given that there are $K$ hyperedges, the total complexity is $O(NKd)$. For the Weighter, the computation is from the dim reduction operation and the linear head. The dim reduction operation is an MLP. In this work, the hidden dimensions are a fraction of the original feature dimension. Therefore, the complexity of the dim-reduction MLP is $O(d^{2})$. The linear head only contributes $O(d)$, which is neglectable. As a result, the computational complexity of the whole model is $O(NKd+d^{2})=O(N^{2}K)$ since the feature dimension is equal to the number of regions (5.1). This complexity is just at the same scale as that of MLPs even though we are addressing a more challenging task: identifying high-order relationships in an exponential space.
5 Experiments
In this section, we conducted experiments to validate the quality of the learned hyperedges regarding the cognition phenotype outcome. Furthermore, we conducted ablation studies to validate the key components in our model. We also analyzed our results both quantitatively and qualitatively.
5.1 Experiment Settings
Evaluation Metric
To evaluate the quality of hyperedges obtained by HyBRiD, we use CPM (Shen et al., 2017), a standard model that could evaluate the correlation between connectivity and the prediction target, due to its high impact in the community. In the original implementation of CPM, weights of pairwise edges are obtained by Pearson correlation between nodes. These weights, as pairwise connectivity, are fed into the CPM. CPM will output a metric that measures the overall correlation between edges and the prediction target, which can be considered as a measure of edge qualities. This process is formulated as
$r^{\prime}=\operatorname{CPM}(\bm{w}_{p},Y)$ (12)
where $\bm{w}_{p}\in\mathbb{R}^{K_{p}}$ denotes the pairwise edge weights and $K_{p}$ is the total number of pairwise edges. $r^{\prime}$ is a metric that measures the quality of weights based on positive and negative correlations to the cognition score. To evaluate the quality of the learned weights for our model and baselines, we concatenate the learned weights $\bm{w}_{h}\in\mathbb{R}^{K_{h}}$ with the pairwise ones together into the CPM, and thus adjust Equation 12 to
$r=\operatorname{CPM}([\bm{w}_{p}\|\bm{w}_{h}],Y)$ (13)
[…]
Datasets
We use functional brain imaging data from the first release of the Adolescent Brain Cognitive Development (ABCD) study, collected from $11,875$ children aged between 9 to 10 years old (Casey et al., 2018). The functional MRI (fMRI) data is collected from children when they were resting and when they performed three emotional and cognitive tasks. The fMRI data is processed using BioImage Suite (Joshi et al., 2011) with the standard preprocessing procedures, such as slice time and motion correction, registration to the MNI template (see details in Greene et al. (2018) and Horien et al. (2019)). We delete scans with more than 0.10 mm mean frame-to-frame displacement. We use the ABCD imaging data collected from the baseline (release 2.0) as well as the 2-year follow-up (release 3.0). For each time point, we included children subjected to four conditions: the resting state where the brain is not engaged in any activity (Rest), the emotional n-back task (EN-back), the Stop Signal task (SST), and the Monetary Incentive Delay (MID) task. In conclusion, we obtain $8$ datasets from $2$ timepoints and $4$ conditions (Rest, SST, EN-back, MID). Statistics of each dataset are summarized in Appendix C.
For the prediction target, we consider fluid intelligence as our label. Fluid intelligence reflects the general mental ability and plays a fundamental role in various cognitive functions. Recent literature has seen success in predicting intelligence based on pairwise predictive modeling (Dubois et al., 2018a; b), and we aim to improve the prediction accuracy of the current methods.
Data preprocessing
The raw fMRI data of an instance is represented in four dimensions (3 spatial dimensions + 1 temporal dimension), which can be imagined as a temporal sequence of 3D images. First, brain images are parceled into regions (or nodes) using the AAL3v1 atlas (Rolls et al., 2020). Following previous works (Kan et al., 2022b; Li et al., 2021; Thomas et al., 2022), each region’s time series is obtained by averaging all voxels in that region. Consistent with previous connectivity-based methods (Li et al., 2021; Kan et al., 2022b; Ktena et al., 2018; Said et al., 2023), for each region, we use its Pearson correlation coefficients to all regions as its features. We randomly split the data into train, validation, and test sets in a stratified fashion. The split ratio is 8:1:1.
Baselines
We compare our method with $3$ classes of baselines: 1) standard method, which is exactly the classical method that predicts outcomes based on pairwise edges (Shen et al., 2017; Dadi et al., 2019; Wang et al., 2021). The comparison with standard methods shows whether the high-order connectivity has its advantage over the classical pairwise one or not. 2) hypergraph construction methods. We consider $k$NN (Huang et al., 2009), $l_{1}$ hypergraph (Wang et al., 2015), and $l_{2}$ hypergraph (Jin et al., 2019). 3) Connectivity-based cognition prediction methods, which are state-of-the-art predictive models based on brain connectivity. We consider BrainNetGNN (Mahmood et al., 2021), BrainGNN (Li et al., 2021), and BrainNetTF (Kan et al., 2022b). BrainGB (Cui et al., 2022a) is a study of different brain graph neural network designs and we include its best design as a baseline. Note that since these models are unable to identify hyperedge structures of brain regions, we input their last layer embeddings (each entry as a weight) into the CPM model. Note that our weights $\bm{w}$ are also last layer embeddings in HyBRiD.
Implementation & Training Details
Due to the data scarcity, training on individual datasets would result in serious overfitting. To mitigate this, we train our model as well as baselines using all eight datasets together. Details of the overfitting issue and our parameter-sharing strategy are further discussed in Appendix E.1. See Appendix D for other implementation and training details such as hyperparameter choices and software/hardware specifications.
5.2 Quality of Hyperedges
We report $r$ values by CPM in Table 1. As we can see, HyBRiD outperforms the state-of-the-art predictive models on $7$ datasets, with an average improvement of $12.1\%$. Moreover, HyBRiD outperforms all other deep learning baselines in efficiency, with $87\%$ faster than the next fastest one (BrainNetTF). Refer to Appendix F for more runtime details.
Type Model SST 1 EN-back 1 MID 1 Rest 1 SST 2 EN-back 2 MID 2 Rest 2
Standard pairwise $0.113$ $0.218$ $0.099$ $0.164$ $0.201$ $0.322$ $0.299$ $0.289$
Hypergraph Construction kNN $0.115$ $0.268$ $0.168$ $0.127$ $0.257$ $0.266$ $0.238$ $0.315$
$l_{1}$ hypergraph $0.099$ $0.223$ $0.125$ $0.126$ $0.145$ $0.295$ $0.242$ $0.259$
$l_{2}$ hypergraph $0.096_{\pm 0.002}$ $0.197_{\pm 0.003}$ $0.118_{\pm 0.003}$ $0.157_{\pm 0.016}$ $0.203_{\pm 0.005}$ $0.272_{\pm 0.004}$ $0.289_{\pm 0.011}$ $0.307_{\pm 0.006}$
Connectivity based Prediction BrainNetGNN $0.227_{\pm 0.060}$ $0.287_{\pm 0.043}$ $0.266_{\pm 0.046}$ $0.221_{\pm 0.040}$ $0.468_{\pm 0.058}$ $0.480_{\pm 0.068}$ $0.506_{\pm 0.057}$ $0.453_{\pm 0.028}$
BrainGB $0.190_{\pm 0.073}$ $0.214_{\pm 0.051}$ $0.265_{\pm 0.048}$ $0.176_{\pm 0.066}$ $0.447_{\pm 0.089}$ $0.483_{\pm 0.077}$ $0.458_{\pm 0.064}$ $0.432_{\pm 0.076}$
BrainGNN $0.262_{\pm 0.030}$ $0.235_{\pm 0.032}$ $0.260_{\pm 0.049}$ $0.185_{\pm 0.058}$ $0.455_{\pm 0.028}$ $0.391_{\pm 0.077}$ $0.445_{\pm 0.078}$ $0.368_{\pm 0.041}$
BrainNetTF $\underline{0.327_{\pm 0.084}}$ $\underline{0.338_{\pm 0.056}}$ $\underline{0.370_{\pm 0.098}}$ $\mathbf{0.334_{\pm 0.084}}$ $\underline{0.633_{\pm 0.178}}$ $\underline{0.631_{\pm 0.142}}$ $\underline{0.629_{\pm 0.123}}$ $\underline{0.588_{\pm 0.138}}$
Ours HyBRiD $\mathbf{0.361_{\pm 0.058}}$ $\mathbf{0.348_{\pm 0.061}}$ $\mathbf{0.386_{\pm 0.060}}$ $\underline{0.223_{\pm 0.056}}$ $\mathbf{0.738_{\pm 0.054}}$ $\mathbf{0.714_{\pm 0.037}}$ $\mathbf{0.816_{\pm 0.053}}$ $\mathbf{0.730_{\pm 0.049}}$
Table 1: $r$ values of our hyperedges compared to baselines. Results are averaged over 10 runs. Deterministic methods do not have standard deviations.
Ablation Studies
[…]
Model SST 1 EN-back 1 MID 1 Rest 1 SST 2 EN-back 2 MID 2 Rest 2
HyBRiD $\mathbf{0.361_{\pm 0.058}}$ $\mathbf{0.348_{\pm 0.061}}$ $\mathbf{0.386_{\pm 0.060}}$ $\underline{0.223_{\pm 0.056}}$ $\mathbf{0.738_{\pm 0.054}}$ $\mathbf{0.714_{\pm 0.037}}$ $\mathbf{0.816_{\pm 0.053}}$ $\mathbf{0.730_{\pm 0.049}}$
[…]
Table 2: Ablation studies on the masking mechanism. Results are averaged over 10 runs.
5.3 Further Analysis
Hyperedge degree distribution
We plot the hyperedge degree distribution in Fig.3(a). We find there are two distinct clusters in the figure. The first cluster is hyperedges with degree $\leq 5$. $1$-degree and $2$-degree hyperedges are special cases of our method: $1$-degree hyperedges are individual nodes, which imply the contribution of individual regions to the cognition. $2$-degree hyperedges reveal the importance of traditional pairwise connectivity. The other cluster concentrates around degree $25$, which implies the importance of relationships of multiple regions.
Hyperedges with higher degree are more significant
CPM conducts a significance test on pairwise edges and hyperedges internally based on a linear regression model, and thus we can obtain a P-value for each hyperedge from the significance test. We define the significance of a hyperedge as $1-P_{v}\in[0,1]$ where $P_{v}$ is the P-value of that hyperedge.
Figure 4: Hyperedge profiles. (a) Hyperedge degree distribution of learned hyperedges. (b) Correlation between hyperedge degree and significance. (c) Comparison between the number hyperedges and pairwise edges under different significance thresholds. The total number of hyperedges is $32$. And the total number of pairwise edges is $26,896$.
The relationship between hyperedge degree and its significance is shown in Fig 3(b). In this figure, we find a strong positive correlation between a hyperedge’s degree and its significance, which indicates that interactions of multiple brain regions play more important roles in cognition than pairwise or individual ones. It is also worth mentioning that there is a turning point around degree $5$, which corresponds to the valley around $5$ in Figure 3(a).
Comparison with pairwise edges
To compare the significance in cognition between pairwise edges and learned hyperedges, we plot the number of remaining edges under different thresholds in Figure 3(c). We find out that the learned hyperedges are much more significant than pairwise ones. Also note that even if we set the threshold to an extremely strict value ($1\times 10^{-8}$), there are still $60\%$ hyperedges considered significant. This evidence shows that our high-order relationships are much more significant than the traditional pairwise connectivity, which implies relationships involving multiple brain regions could be much more essential in cognition.
Region importance
Figure 5: Visualization of the frequency of each region under different conditions.
To better understand the roles of each brain region in cognition under different conditions, we studied the frequency at which each region appears in a hyperedge out of all identified hyperedges. The frequency, which can be considered as a measure of region importance, is visualized in Figure 5. Visual regions (Fusiform, Cuneus, Calcarine) are especially active due to the intensive visual demands in all three conditions. We found that the Frontal_Inf_Oper and ACC_pre, recognized for their participation in response inhibition (Pornpattananangkul et al., 2016), frequently appear in the SST task. This aligns with what the SST task was designed to test. Interestingly, of the three conditions (SST, EN-back, MID), only EN-back prominently involves the Amygdala, a key region for emotion processing. This makes sense as EN-back is the only condition related to emotion. More visualizations and interpretations on the resting-state region importance and individual hyperedges can be found in Appendix G.
6 Conclusion
In this work, we propose HyBRiD for identifying maximally informative yet minimally redundant (MIMR) high-order relationships of brain regions. To effectively optimize our model, we further proposed a novel information bottleneck framework with theoretical guarantees. Our method outperforms state-of-the-art models of hypergraph construction and connectivity-based prediction. The result analysis shows the effectiveness of our model. We expect such advancements could benefit clinical studies, providing insights into neurological disorders, and offering improved diagnostic tools in neurology.
Limitations
HyBRiD only considers static high-order relations. Given that ABCD tasks are dynamic, including temporal changes and interactions, it will be interesting to study the evolution of these high-order relationships.
Appendix A Proof of the Upper Bound
In this section, we prove the upper bound of $I(H;X)$ in multi-head drop-bottleneck in Equation 10.
$y$$x$$z$
Figure 6: The graphical model of random variables $X$, $Y$ and $Z$.
Lemma 1.
Given random variables $X$, $Y$ and $Z$. Their relationships are described in the graphical model illustrated in Figure 6. We have
$I(X;Y|Z)\leq I(X;Y)$ (14)
Proof.
$I(X;Y|Z)-I(X;Y)$ $=\int p(x,y,z)\log\frac{p(z)p(x,y,z)}{p(x,z)p(y,z)}\mathrm{d}x\mathrm{d}z\mathrm{d}y$ (15)
$-\int p(x,y)\log\frac{p(x,y)}{p(x)p(y)}\mathrm{d}x\mathrm{d}y$
$=\int p(x,y,z)\log\frac{p(z)p(x,y,z)}{p(x,z)p(y,z)}\mathrm{d}x\mathrm{d}z\mathrm{d}y$
$-\int p(x,y,z)\log\frac{p(x,y)}{p(x)p(y)}\mathrm{d}x\mathrm{d}y\mathrm{d}z$
$=\int p(x,y,z)\log\frac{p(z)p(x,y,z)p(x)p(y)}{p(x,z)p(y,z)p(x,y)}\mathrm{d}x\mathrm{d}z\mathrm{d}y$
$=\int p(x,y,z)\log\frac{p(z)p(x,z|y)p(y)p(x)p(y)}{p(x,z)p(y,z)p(x,y)}\mathrm{d}x\mathrm{d}z\mathrm{d}y$
$=\int p(x,y,z)\log\frac{p(z)p(x|y)p(z|y)p(y)p(x)p(y)}{p(x,z)p(y,z)p(x,y)}\mathrm{d}x\mathrm{d}z\mathrm{d}y$
$=\int p(x,y,z)\log\frac{p(x)p(z)}{p(x,z)}\mathrm{d}x\mathrm{d}z\mathrm{d}y$
$=\int p(x,z)\log\frac{p(x)p(z)}{p(x,z)}\mathrm{d}x\mathrm{d}z$
$=-I(X;Z)\leq 0$
which finishes the proof.
Corollary 1.1.
Given the same graphical model 6, we have
$I(X,Z;Y)\leq I(X;Y)+I(Z;Y)$ (16)
Proof.
Using the chain rule of mutual information, we obtain
$I(X,Z;Y)=I(X;Y|Z)+I(Z;Y)$ (17)
According to Lemma 1, we have $I(X;Y|Z)\leq I(X;Y)$, which finishes the proof.
Theorem 2.
For random variables in Equation 10, we have
$I(H;X)\leq\sum_{k=1}^{K}I(\bm{h}^{k};X)$ (18)
[…]
Figure 7: The graphical model of random variables $\bm{h}^{k}$ and $X$
Proof.
According to the definitions of $X$ and $H$, which are described in Section 4, we can draw a graphical model of them in Figure 7. Define a new random variable $\bm{h}^{k_{1}:k_{2}}=[\bm{h}^{k_{1}},\bm{h}^{k_{1}+1},\cdots,\bm{h}^{k_{2}-1},\bm{h}^{k_{2}}]$, which is a concatenation from $\bm{h}^{k_{1}}$ to $\bm{h}^{k_{2}}$. According to Corollary 1.1 we have
$I(H;X)$ $\leq I(\bm{h}^{1},X)+I(\bm{h}^{2:K},X)$ (19)
$\leq I(\bm{h}^{1},X)+I(\bm{h}^{2},X)+I(\bm{h}^{3:K},X)$
$\leq I(\bm{h}^{1},X)+I(\bm{h}^{2},X)+I(\bm{h}^{3},X)+\cdots$
$\leq\sum_{k=1}^{K}I(\bm{h}^{k};X)$
Theorem 3.
(proposition 1)
$I(H;X)\leq\sum_{k=1}^{K}\sum_{i=1}^{N}I(\bm{h}^{k}_{i};X_{i})=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i})$ (20)
Proof.
Given Theorem 2. It suffices to prove
$I(\bm{h}^{k};X)\leq\sum_{i=1}^{N}I(\bm{h}^{k}_{i};X_{i})=\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i}),\quad\forall 1\leq k\leq K$ (21)
And this is exactly the conclusion in Kim et al. (2021) if we consider $\bm{h}^{k}$ and $X$ as $X$ and $Z$ respectively in their paper. ∎
Appendix B Connectome-Based Predictive Modeling
Shen et al. (2017); Finn et al. (2015)) has shown tremendous promise in recent years in detecting imaging biomarkers. (Rosenberg et al., 2015; Dubois et al., 2018a; Rosenberg et al., 2020; 2016). Such a model, based chiefly on functional MRI data, can measure the significance of the input edge weights, which is revealed by a correlation coefficient that reflects the correlation between the edge weights and the neurological outcomes. One could expect a large correlation coefficient to indicate the high quality of edge weights. We utilize the CPM as an evaluation model to evaluate the quality of our learned hyperedges. Here is a pipeline overview of the CPM process:
  1. 1.
    Connectivity Calculation: For each subject, compute the Pearson correlation coefficients for each possible pair of brain regions. This is based on the fMRI series of those regions.
  2. 2.
    Edge Significance: Calculate the correlation between each brain connectivity edge and the outcome of interest (e.g., cognition scores) across all subjects. The correlation of an edge indicates its significance.
  3. 3.
    Edge Selection: Identify significant connectivity edges. These are the edges where the correlation values are greater than a predetermined significance threshold.
  4. 4.
    Weight Summation: For each subject, sum the weights of the significant edges identified in the previous step to derive a single summary score (scalar).
  5. 5.
    Model Fitting: Fit a linear model that predicts the neurological outcomes based on the summed weights, where each subject is a sample.
  6. 6.
    Model Evaluation: Across all subjects, calculate the correlation of predicted values and the neurological outcomes. Note that this correlation coefficient is equivalent to the one between the summed weights and the outcomes, and is exactly the metric $r$ we use to evaluate our hyperedges in Equation 13.
Since positive edges and negative edges will cancel out with each other when being summed, we adopt the combining strategy in Boyle et al. (2023).
CPM measures the quality of edge weights
According to step 6, the evaluation of the predictive model could be measured by the correlation between predicted and ground-truth outcomes Shen et al. (2017). Since CPM is a linear model that predicts the outcome based on the sum of significant edge weights, the correlation is equal to the correlation between the sum and ground-truth outcomes (which is exactly the r in Eq. 13). Hence, one can expect a larger correlation if the edge weights are more correlated (and thus are more predictive).
Significance of edges in CPM
In step 2, CPM obtains a correlation coefficient $r^{k}$ for each edge weight $w^{k}$ and the cognition score $Y$ across all subjects. Consider a classical hypothesis test $H_{0}:r^{k}=0,H_{1}:r^{k}\neq 0$. Assume $w^{k}$ and $Y$ are drawn from independent normal distribution (corresponds to $H_{0}$), the probability density function of correlation coefficient $r^{k}$ is
$f(r^{k})=\frac{\left(1-{r^{k}}^{2}\right)^{n/2-2}}{\mathrm{~{}B}\left(\frac{1}{2},\frac{n}{2}-1\right)},$
where $n$ is the number of samples and $B$ is the beta function. Based on the distribution, we obtain a $P$-value for the $k$-th edge, which is used to measure the significance of the edge.
Appendix C Dataset details
The statistics of the number of instances and the time series length are summarized in Table 3.
Dataset Rest 1 SST 1 EN-back 1 MID 1 Rest 2 SST 2 EN-back 2 MID 2
#(instances) 1676 1673 1678 1678 1949 1053 1044 1062
length of time series 375 437 362 403 375 437 362 403
Table 3: Statistics of 8 datasets.
Appendix D Training Details
Hardware
We train our model on a machine with a Intel Xeon Gold 6326 CPU and RTX A5000 GPUs.
Software
See Table 4 for software we used and the versions.
software version
python 3.8.13
pytorch 1.11.0
cudatoolkit 11.3
numpy 1.23.3
ai2-tango 1.2.0
nibabel 4.0.2
Table 4: Software versions
Hyperparameter choices
The hyperparameters selection is shown in Table 5. Some crucial hyperparameters ablation experiments can be found in Appendix E.
notation meaning value
$lr$ learning rate $1\times 10^{-3}$
$K$ number of hyperedges $32$
$\beta$ trade-off coefficients information bottleneck $0.2$
$[h_{1},h_{2},h_{3}]$ hidden sizes of the dim reduction MLP $[32,8,1]$
$B$ batch size $64$
Table 5: Hyperparameter choices
Appendix E More Ablation Studies
E.1 Parameters Sharing
As previously discussed, the scarcity of fMRI data can result in serious over-fitting. In our experiments, we observed that when the model is trained on each dataset individually, the validation set loss remains stagnant, which indicates over-fitting. To address this, we examined the model’s performance under various parameter-sharing strategies. Specifically, we have three learnable components: the shallow embeddings in Equation 3, the dim reduction MLP in 6 and the LinearHead $\mathcal{F}_{l}$. These components can be either shared or condition-specific. Results of different parameter-sharing strategy is shown in Table 6.
Strategy ID ShallowEmbeddings DimReduction LinearHead Average Score
1 $0.337$
2 $0.335$
3 $0.508$
4 $\underline{0.539}$
5 $0.351$
6 $0.384$
7 $0.506$
8 $\mathbf{0.558}$
Table 6: The table summarizes the model’s average performance under various parameter-sharing strategies. The symbol ✓ indicates a component being condition-specific, while an empty cell implies a component being shared across conditions.
Note that strategy 1 is equivalent to training models separately for each condition. From the table, we find that the DimReduction component plays the most significant role. Making DimReduction condition-specific enhances the average performance by $10\%-20\%$. Strategy 8 is the best. However, domain experts may expect different structures under different conditions. Therefore, we still adopt condition-specific ShallowEmbeddings (i.e. strategy 4), given that their scores are really close. We argue that the fact that strategy 4 is not better than strategy 8 is because our prediction target is fluid intelligence, which is a general measurement of cognition and thus tends to be condition-agnostic. We anticipate a potential shift in results if we switch to a more condition-specific cognition phenotype outcome.
E.2 Choices of number of hyperedges $K$
As explained in Section 4, we use $K$ heads for $K$ hyperedges. We study the correlation between the $r$ value and the number of hyperedges on three datasets:
(a) The performance with increasing the number of hyperedges through $2$, $7$, $12$, $17$, $22$, $27$, $32$, $37$, $52$, $57$, $62$ on three datasets.
(b) Performance with increasing $\beta$ from $0.1$ to $1.0$ with step $0.1$ on all datasets.
From Figure 7(a), we find that the overall performance increases dramatically before $K=17$, but becomes stable and close to saturation after $K=32$. To improve the efficiency while ensuring the performance, we choose $K=32$.
E.3 Choices of the trade-off coefficient $\beta$
In our optimization objective 8, $\beta$ acts as a trade-off parameter, which is a non-negative scalar that determines the weight given to the second term relative to the first. To study its fluence to the performance, we plot the model performances on all datasets under different $\beta$ in Figure 7(b). We can see performances on $3$ datasets (Rest 1, SST 2) consistently decrease when $\beta$ increases. However, on other $5$ datasets (SST 1, MID 1, Rest 2, nBack 2, MID 2), we can observe a peak at $\beta=0.2$. Accordingly, we adopt $\beta=0.2$.
Appendix F Runtime Comparison
Figure 8 summarizes the per-batch training time of all deep learning models. We find that HyBRiD is the most efficient one, with $87\%$ faster than the second one (BrainNetTF) and at least $1255\%$ faster than the GNN-based ones.
Figure 8: Training time per batch of all deep learning models.
Appendix G More Visualizations
Hyperedge case study
We visualize the most significant hyperedge under the EN-back condition. We observe a coordinated interaction of numerous brain regions, each fulfilling specific roles. Notably, some of these regions serve multi-functional purposes:
  • Memory Processing ParaHippocampal_L, Temporal Mid: Essential for memory encoding and retrieval, these regions are integral to the EN-back task, facilitating the recall of previously viewed images.
  • Emotional processing Amygdala_R: The amygdala is crucial for the processing of emotions, such as fear and pleasure. Since the EN-back task involves emotional stimuli, it is reasonable that the region is connected by the hyperedge.
  • Visual Processing: Calcarine_R, Lingual_L, Fusiform_L. These regions are responsible for visual perception and some of them are related to complex visual contents like symbols and human faces, which were presented during the task.
  • Sensory SupraMarginal_L: It is responsible for interpreting tactile sensors and perceiving limbs location. Its involvement is likely due to the requirement for participants to engage in specific physical actions, such as pressing buttons, during the task. Temporal Mid: It functions in multi-modal sensory integration.
  • Motor Control Cerebellum: It is primarily responsible for muscle control. Caudate: It plays a crucial role in motor processes. Its involvement is likely attributed to participants engaging in physical actions, like pressing buttons.
  • Cognitive Control ACC_pre_L: In the EN-back task, this region is likely crucial for maintaining focus, error detection and correction, conflict management in working memory, and modulating emotional responses to the task’s demands.
Figure 9: Visualization of the most significant hyperedge in EN-back condition
Resting-state brain region importance
We visualize the region importance of the resting state in Figure 10. Different from task states, where specific brain regions are activated in response to particular tasks, brain activities during resting states, are not driven by external tasks, leading to more diffuse and less predictable patterns of activation. This makes it harder to pinpoint specific interactions or functions.
Figure 10: Region importance of resting state.
body>

### v2

@@
 Abstract
-Discovering reliable and informative interactions among brain regions from functional magnetic resonance imaging (fMRI) signals is essential in neuroscientific predictions of cognition.
+Discovering reliable and informative relationships among brain regions from functional magnetic resonance imaging (fMRI) signals is essential in phenotypic predictions.
 Most of the current methods fail to accurately characterize those interactions because they only focus on pairwise connections and overlook the high-order relationships of brain regions.
-We delve into this problem and argue that these high-order relationships should be maximally informative and minimally redundant (MIMR).
-However, identifying such high-order relationships is challenging and highly under-explored.
-Methods that can be tailored to our context are also non-existent.
-In response to this gap, we propose a novel method named HyBRiD that aims to extract MIMR high-order relationships from fMRI data.
-HyBRiD employs a Constructor to identify hyperedge structures, and a Weighter to compute a weight for each hyperedge.
+We propose that these high-order relationships should be maximally informative and minimally redundant (MIMR).
+However, identifying such high-order relationships is challenging and under-explored due to the exponential search space and the absence of a tractable objective.
+In response to this gap, we propose a novel method named HyBRiD which aims to extract MIMR high-order relationships from fMRI data.
+HyBRiD employs a Constructor to identify hyperedge structures, and a Weighter to compute a weight for each hyperedge, which avoids searching in exponential space.
 HyBRiD achieves the MIMR objective through an innovative information bottleneck framework named multi-head drop-bottleneck with theoretical guarantees.
 Our comprehensive experiments demonstrate the effectiveness of our model.
-Our model outperforms the state-of-the-art predictive model by an average of $12.1\%$, regarding the quality of hyperedges measured by CPM, a standard protocol for studying brain connections.
+Our model outperforms the state-of-the-art predictive model by an average of $11.2\%$, regarding the quality of hyperedges measured by CPM, a standard protocol for studying brain connections.
+Machine Learning, ICML
 1 Introduction
-Discovering reliable and informative relations among brain regions using fMRI signals is crucial for predicting human cognition in and understanding brain functions (Kucian et al., 2008; 2006; Li et al., 2015b; a; Satterthwaite et al., 2015; Wang et al., 2016).
-However, despite the clear multiplexity of the brain’s involvement in cognition (Logue & Gould, 2014; Barrasso-Catanzaro & Eslinger, 2016; Knauff & Wolf, 2010; Reineberg et al., 2022), the current imaging biomarker detection methods (Shen et al., 2017; Gao et al., 2019; Li et al., 2021) focus only on the contributing roles of the pairwise connectivity edges.
+Discovering relations among brain regions toward a specific phenotypic outcome from fMRI signals has been a crucial area in neuroimaging research.
+Reliable and informative relations help neuroscientists and clinical professionals to better understand brain functions, and thus improve clinical diagnosis and treatments (Kucian et al., 2008, 2006; Li et al., 2015b, a; Satterthwaite et al., 2015; Wang et al., 2016).
+However, despite the clear multiplexity of the brain’s involvement in cognition (Logue & Gould, 2014; Barrasso-Catanzaro & Eslinger, 2016; Knauff & Wolf, 2010; Reineberg et al., 2022), current imaging biomarker detection methods (Shen et al., 2017; Gao et al., 2019; Li et al., 2021) focus only on the contributing roles of the pairwise connectivity edges.
 In contrast, most brain functions involve distributed patterns of interactions among multiple regions (Semedo et al., 2019).
@@
 Pairwise connectivity, without considering the brain’s high-order structure, can lead to inconsistent findings across studies with low prediction performance.
-Although there are few works (Zu et al., 2016; Xiao et al., 2019; Li et al., 2022) working on discovering the high-order relationships of brain regions, they cannot effectively extract meaningful patterns towards a neurological outcome since 1) the discovering process of high-order relationships is not guided by cognitive targets; 2) the weak expressiveness of traditional machine learning methods (e.g.
-lasso regression) hinders the accurate modeling of brain region interactions (Cao et al., 2022; Richards et al., 2019).
-In this paper, we aim to identify high-order relationships that are informative toward a neurological outcome.
+Although recently there have been a few works (Zu et al., 2016; Xiao et al., 2019; Li et al., 2022) working on discovering the high-order relationships of brain regions, they are unable to effectively extract meaningful patterns.
+This is because these methods usually first identify some candidates of high-order relationships and then perform feature selection.
+The candidates are obtained by enumerating only a small portion of all possible relations, or by clustering methods that are unrelated to the target.
+As a result, most informative high-order relations are likely excluded at the first stage.
+This inspires us to solve the problem in an end-to-end manner through a more expressive model.
+In this paper, we aim to identify high-order relationships that are informative towards a phenotypic outcome, such as a cognition score.
 However, unlike pairwise relations, the number of possible high-order relations is exponential.
 To identify the most informative ones from the exponential space, we propose our objective: maximally informative and minimally redundant (MIMR).
-That is to say, we aim to maximize the information contained in the high-order relationships towards a neurological outcome (informativeness) while diminishing the participation of unrelated brain regions (redundant).
-Such a criterion, on the one hand, ensures the predictive performance of these high-order relationships; on the other hand, it endows the model with the capacity to identify more succinct and interpretable structures (Yu et al., 2020; Miao et al., 2022a; b; Chen & Ying, 2023).
-A formal definition of the MIMR criterion could be found in Equation 8 from an information bottleneck view.
-Figure 1: We identify high-order relationships of brain regions, where hyperedges possess strong relevance to cognition (maximal informativeness).
+That is to say, we maximize the information contained in the high-order relationships towards a neurological outcome (informativeness) while diminishing the participation of unrelated brain regions (redundant).
+Such a criterion, on the one hand, ensures the predictive performance of these high-order relationships; on the other hand, it endows the model with the capacity to identify more succinct and interpretable structures (Yu et al., 2020; Miao et al., 2022a, b; Chen & Ying, 2023).
+A formal definition of the MIMR criterion could be found in Equation 8 from an information bottleneck point of view.
+Figure 1: We identify high-order relationships of brain regions, where hyperedge structures and weights possess strong relevance to cognition (maximal informativeness).
 Meanwhile, they contain the least irrelevant information.
-We formulate high-order relationships as hyperedges in a hypergraph, where regions are treated as nodes.
+We formulate high-order relationships as weighted hyperedges in a hypergraph, where regions are treated as nodes.
 Unlike a traditional graph where edges connect only two nodes, a hypergraph allows edges, known as hyperedges, to connect any number of nodes.
-The hypergraph should be weighted, and the weights of hyperedges are considered as strengths of high-order relationships, which contain the information relevant to cognition (Figure 1).
-However, current methods for hypergraph construction, which mostly are based on neighbor distances and neighbor reconstruction (Wang et al., 2015; Liu et al., 2017; Jin et al., 2019; Huang et al., 2009), are unsuitable in our context for several reasons: 1) they are unable to learn MIMR hyperedges due to the absence of a tractable objective for learning such hyperedges.
+The hypergraph should be weighted, and the weights of hyperedges are considered as strengths of high-order relationships, which contain the information relevant to the outcome (Figure 1).
+However, current methods for hypergraph construction, which are mostly based on neighbor distances and neighbor reconstruction (Wang et al., 2015; Liu et al., 2017; Jin et al., 2019; Huang et al., 2009), are unsuitable in our context for several reasons: 1) they are unable to learn MIMR hyperedges due to the absence of a tractable objective for learning such hyperedges.
 2) they fall short of learning consistent structures across subjects, which contradicts the belief that the cognitive function mechanism of healthy human beings should be similar (Wang et al., 2023).
 3) the number of hyperedges is restricted to the number of nodes, which may lead to sub-optimal performance.
-Furthermore, although information bottleneck (IB) is a prevailing solution to learn MIMR representations in deep learning (Kim et al., 2021; Alemi et al., 2016; Luo et al., 2019), existing IB methods focus on extracting compressed representations of inputs instead of identifying underlying structures such as hypergraphs.
+Furthermore, although information bottleneck (IB) has been a prevailing solution to learn MIMR representations in deep learning (Kim et al., 2021; Alemi et al., 2016; Luo et al., 2019), existing IB methods focus on extracting compressed representations of inputs instead of identifying underlying structures such as hypergraphs.
 Harnessing the IB framework for identifying hypergraphs necessitates both architectural innovations and theoretical derivations.
@@
 To advance the IB principle for hyperedge identification, we further propose multi-head drop-bottleneck and derive its optimization objective.
-HyBRiD gets rid of searching in an exponential space through learning masks to identify hyperedges, which guarantees computational efficiency.
+HyBRiD avoids searching in an exponential space through learning masks to identify hyperedges, which guarantees efficiency.
 Its feature-agnostic masking mechanism design ensures HyBRiD to learn consistent structures across subjects.
-Moreover, the model is equipped with a number of parallel heads, and each head is dedicated to a hyperedge.
+Moreover, the model is equipped with a number of parallel heads, each of which is dedicated to a hyperedge.
 Through this, HyBRiD is able to identify any number of hyperedges, depending on how many heads it is equipped with.
 Additionally, the proposed multi-head drop-bottleneck theoretically guarantees the maximal informativeness and minimal redundancy of the identified hyperedges.
-We evaluate our methods on $8$ datasets of human subjects under different conditions and at different participation times.
-We quantitatively evaluate our approach by a commonly used protocol for studying brain connections, CPM (Shen et al., 2017) (Appendix B), and show that our model outperforms the state-of-the-art deep learning models by an average of $12.1\%$ on a comprehensive benchmark.
+We evaluate our methods on the open-source ABIDE dataset and the restricted ABCD dataset.
+We quantitatively evaluate our approach by a commonly used protocol for studying brain connections, CPM (Shen et al., 2017) (Appendix B), and show that our model outperforms the state-of-the-art deep learning models by an average of $11.2\%$ on a comprehensive benchmark.
 Our post-hoc analysis demonstrates that hyperedges of higher degrees are considered more significant, which indicates the significance of high-order relationships in human brains.
@@
 Input
-We consider ABCD, which is a widely used benchmark in fMRI studies (5.1).
-This dataset is a collection of instances, where each instance corresponds to an individual human subject.
-An instance is represented by the pair $(X,Y)$.
-$X\in\mathbb{R}^{N\times d}$ represents the features for each subject, where $N$ is the number of brain regions and $d$ is the number of feature dimensions.
-The features are obtained by connectivity, consistent with previous works (5.1).
-$Y\in\mathbb{R}$ denotes the prediction target, such as fluid intelligence.
-Section 5 elaborates more details about datasets and preprocessing procedures.
-For how to obtain features of regions from raw data, see Paragraph 5.1 for more details.
+Our dataset is a collection of human subject’s features and their phenotypic outcomes, which is represented by the pair $(X,Y)$ for each subject.
+$X\in\mathbb{R}^{N\times d}$ represents the features for each subject, where $N$ is the number of brain regions and $d$ is the feature size.
+Consistent with previous works (Kan et al., 2022b; Li et al., 2021), the features are Pearson correlations derived from fMRI time series.
+$Y\in\mathbb{R}$ denotes the phenotypic outcome, such as the intelligent quotient.
+Section 5 and Appendix D will elaborate more details about datasets and preprocessing procedures.
 Goal
 Based on the input $X$, HyBRiD aims to learn a weighted hypergraph of the brain, where regions are nodes.
-To achieve this, HyBRiD identifies a collection of hyperedges $H=(\bm{h}^{1},\bm{h}^{2},\cdots,\bm{h}^{K})$, and assigns weights $\bm{w}=[w^{1},w^{2},\cdots,w^{K}]^{T}$ for all hyperedges.
-These hyperedges and their weights, which represent strengths of hyperedges, are expected to be the most informative (i.e.
-relevant to the target $Y$) yet the least redundant.
+To achieve this, HyBRiD identifies a collection of hyperedges $H=(\boldsymbol{h}^{1},\boldsymbol{h}^{2},\cdots,\boldsymbol{h}^{K})$, and assigns weights $\boldsymbol{w}=[w^{1},w^{2},\cdots,w^{K}]^{T}$ for all hyperedges.
+These hyperedges and their weights, which represent strengths of hyperedges, are expected to be the most informative towards $Y$ yet the least redundant.
 Representation of Hyperedges
-As mentioned before, we use $H$ to denote the collection of hyperedge structures and $\bm{h}^{k}$ to denote the $k$-th hyperedge.
+As mentioned before, we use $H$ to denote the collection of hyperedge structures and $\boldsymbol{h}^{k}$ to denote the $k$-th hyperedge.
 To associate hyperedges with node memberships, we use the following representation for a hyperedge:
-$\bm{h}^{k}=\bm{m}^{k}\odot X\in\mathbb{R}^{N\times d}$ (1)
-where $\bm{m}^{k}\in\{0,1\}^{N}$ is a mask vector and $\odot$ denotes broadcasting element-wise multiplication. In other words, each $\bm{h}^{k}$ is a random row-zeroed version of $X$.
+$\boldsymbol{h}^{k}=\boldsymbol{m}^{k}\odot X\in\mathbb{R}^{N\times d},$ (1)
+where $\boldsymbol{m}^{k}\in\{0,1\}^{N}$ is a mask vector and $\odot$ denotes broadcasting element-wise multiplication.
+In other words, each $\boldsymbol{h}^{k}$ is a randomly row-zeroed version of $X$.
 3 Related Work
@@
 Zhang et al.
-(2022; 2018) proposed to iteratively refine a noisy hypergraph, which is obtained by the aforementioned methods.
+(2022, 2018) proposed to iteratively refine a noisy hypergraph, which is obtained by the aforementioned methods.
 Therefore, they share the same limitations as the aforementioned methods.
 In addition, these methods are unable to learn MIMR hyperedges due to the absence of a tractable objective.
-Other methods such as attributed-based methods (Huang et al., 2015; Joslyn et al., 2019) are ill-adapted to our context since they require discrete labels or a prior graph topology.
-Different from these methods, we provide a way to learn a consistent hypergraph through a deep-learning model without any prior topology.
+Other methods, such as attributed-based methods (Huang et al., 2015; Joslyn et al., 2019), are ill-adapted to our context since they require discrete labels.
+Different from these methods, we provide a way to learn a consistent hypergraph through a deep-learning model.
 Furthermore, thanks to the proposed multi-head drop-bottleneck, these hyperedges are theoretically ensured MIMR.
 High-Order Relationships in fMRI
-Although some methods are working on high-order relationships in fMRI, they are limited or inconsistent with our MIMR objective.
+Although there are some methods working on high-order relationships in fMRI, they have limitations and are inconsistent with our MIMR objective.
 Xiao et al.
@@
 After that, IB has been widely used in deep learning.
-The applications span areas such as computer vision (Luo et al., 2019; Peng et al., 2018), reinforcement learning (Goyal et al., 2019; Igl et al., 2019), natural language processing (Wang et al., 2020) and graph learning (Yu et al., 2020; 2022; Xu et al., 2021; Wu et al., 2020).
+The applications span areas such as computer vision (Luo et al., 2019; Peng et al., 2018), reinforcement learning (Goyal et al., 2019; Igl et al., 2019), natural language processing (Wang et al., 2020) and graph learning (Yu et al., 2020, 2022; Xu et al., 2021; Wu et al., 2020).
 Unlike these studies that use IB to extract a compressed representation or a select set of features, our approach focuses on identifying the underlying structures of the data.
-Connectivity-based cognition prediction
-Recently, deep learning techniques have been increasingly employed in predicting cognition based on the connectivity of brain regions.
-A substantial portion of these studies (Ahmedt-Aristizabal et al., 2021; Li et al., 2019; Cui et al., 2022b; Kan et al., 2022a; Cui et al., 2022a; Said et al., 2023) model the brain network as a graph, in which regions act as nodes and pairwise correlations form the edges.
+Connectivity-based Phenotypic Prediction
+Recently, deep learning techniques have been increasingly employed in predicting phenotypic outcomes based on the connectivity of brain regions.
+Most works (Ahmedt-Aristizabal et al., 2021; Li et al., 2019; Cui et al., 2022b; Kan et al., 2022a; Cui et al., 2022a; Said et al., 2023) model the brain network as a graph, in which regions act as nodes and pairwise correlations form the edges.
 These methods predominantly utilize Graph Neural Networks (GNNs) to capture the connectivity information for predictions.
@@
 (2022b) proposed to use transformers with a specially designed readout module, leveraging multi-head attention mechanisms to capture pairwise connectivity.
-However, these methods heavily rely on pairwise connectivity and neglect more intricate higher-order relationships.
-This oversight, on the one hand, leads to sub-optimal prediction performances; and on the other hand, prevents domain experts from acquiring insightful neuroscience interpretations, given that human cognition often involves multiple regions.
+However, all of these methods heavily rely on pairwise connectivity and neglect more intricate higher-order relationships.
+This oversight, on the one hand, leads to sub-optimal prediction performances and, on the other hand, prevents domain experts from acquiring insightful neuroscience interpretations, given that brain functions often involves multiple regions.
 4 Method
@@
 After that, the Weighter $\mathcal{F}_{w}$ calculates a weight for each hyperedge.
-Finally, based on all the weights $\bm{w}$, the LinearHead $\mathcal{F}_{l}$ predicts the label $Y$.
+Finally, based on all the weights $\boldsymbol{w}$, the LinearHead $\mathcal{F}_{l}$ predicts the label $Y$.
 An illustration of this pipeline is shown in Figure 2.
@@
 […]
-We will elaborate on all of the details of the architecture below.
+We will elaborate on the details of the architecture below.
 Figure 2: Overview of the HyBRiD pipeline when the total number of hyperedges $K=3$.
@@
 4.1 Learning the Hypergraph by Multi-head Masking
-Given an instance, represented by $X=[X_{1},X_{2},\cdots,X_{N}]^{T}\in\mathbb{R}^{N\times d}$, where $X_{i}\in\mathbb{R^{d}}$ is a column vector representing the features of region $i$.
+Each instance (human subject) is represented by $X=[X_{1},X_{2},\cdots,X_{N}]^{T}\in\mathbb{R}^{N\times d}$, where $X_{i}\in\mathbb{R^{d}}$ is a column vector representing the features of region $i$.
 These regions are nodes in the hypergraph we are going to construct.
-Hyperedges in the hypergraph can be beneficial in the learning task because it is essential to model the interactions between more than two regions.
+Hyperedges in the hypergraph can be beneficial in the learning process below because it is essential to model the relationships between more than two regions.
 Hyperedges Construction
-In this paragraph, we elaborate how the Constructor identifies the hyperedges, i.e. $H=\mathcal{F}_{c}(X)$.
+In this paragraph, we elaborate on how the Constructor identifies the hyperedges, i.e. $H=\mathcal{F}_{c}(X)$.
 Suppose the number of hyperedges is $K$, which is a predefined hyperparameter.
@@
 Each head is responsible for constructing a hyperedge by selecting nodes belonging to that hyperedge.
-Specifically, to construct the $k$-th hyperedge, the Constructor’s $k$-th head outputs a column vector $\bm{m}^{k}\in\{0,1\}^{N}$, where each element in the vector corresponds to a region.
-$\bm{m}^{k}$ $=[\mathbbm{1}(p^{k}_{\theta,1}),\mathbbm{1}(p^{k}_{\theta,2}),\cdots,\mathbbm{1}(p^{k}_{\theta,N})]^{T}\in\{0,1\}^{N}$ (3)
-where $p^{k}_{\theta,i},i=1,2,\cdots,N$ are learnable parameters.
+Specifically, to construct the $k$-th hyperedge, the Constructor’s $k$-th head outputs a column vector $\boldsymbol{m}^{k}\in\{0,1\}^{N}$, where each element in the vector corresponds to a brain region,
+$\boldsymbol{m}^{k}=[\mathbbm{1}(p^{k}_{\theta,1}),\mathbbm{1}(p^{k}_{\theta,2}),\cdots,\mathbbm{1}(p^{k}_{\theta,N})]^{T}\in\{0,1\}^{N},$ (3)
+where $p^{k}_{\theta,i}\in[0,1],i=1,2,\cdots,N$ are learnable probabilities.
 $\mathbbm{1}:[0,1]\mapsto\{0,1\}$ is an indicator function, which is defined as $\mathbbm{1}(x)=1$ if $x>0.5$ and $\mathbbm{1}(x)=0$ if $x\leq 0.5$.
-And $\bm{m}^{k}$ is a column vector corresponding to the $k$-th hyperedge.
-Note that since there is no gradient defined for the indicator operation, we employ the stopgradient technique (Oord et al., 2017; Bengio et al., 2013) to approximate the gradient.
+And $\boldsymbol{m}^{k}$ is a column vector corresponding to the $k$-th hyperedge.
+Note that since there is no gradient defined for the indicator operation, we employ the stop-gradient technique (Oord et al., 2017; Bengio et al., 2013) to approximate it.
 Figure 3: Architecture details of the Constructor and the Weighter when the number of nodes $N=6$ and the number of hyperedges $K=3$.
 At a high level, the Constructor learns the hyperedge structure by masking nodes.
-The Weighter computes the weight of each hyperedge based on the the hyperedge’s member nodes and their features.
-In the vector $\bm{m}^{k}$, 00 indicates nodes that are masked out, and $1$ indicates nodes that are not masked.
+The Weighter computes the weight of each hyperedge based on the hyperedge’s member nodes and their features.
+In the vector $\boldsymbol{m}^{k}$, 00 indicates nodes that are masked out, and $1$ indicates nodes that are not masked.
 Nodes that are not masked are considered to form a hyperedge together.
-We use $\bm{h}^{k}$ to represent the masked version of $X$
-$\bm{h}^{k}$ $=\bm{m}^{k}\odot X$ (4)
-$=[\bm{m}^{k}_{1}X_{1},\bm{m}^{k}_{2}X_{2},\cdots,\bm{m}^{k}_{N}X_{N}]\in\mathbb{R}^{N\times d}$
-where $\odot$ is the broadcast element-wise multiplication. $\bm{m}^{k}_{j}$ is the $j$-th element of the vector $\bm{m}^{k}$.
+We use $\boldsymbol{h}^{k}$ to represent the masked version of $X$
+$\boldsymbol{h}^{k}$ $=\boldsymbol{m}^{k}\odot X$ (4)
+$=[\boldsymbol{m}^{k}_{1}X_{1},\boldsymbol{m}^{k}_{2}X_{2},\cdots,\boldsymbol{m}^{k}_{N}X_{N}]\in\mathbb{R}^{N\times d},$
+where $\odot$ is the broadcast element-wise multiplication. $\boldsymbol{m}^{k}_{j}$ is the $j$-th element of the vector $\boldsymbol{m}^{k}$.
 We obtain $K$ hyperedges for $K$ sets of masks. We use $H$ to denote the collection of all hyperedges.
-$H=(\bm{h}^{1},\bm{h}^{2},\cdots,\bm{h}^{K})$ (5)
+$H=(\boldsymbol{h}^{1},\boldsymbol{h}^{2},\cdots,\boldsymbol{h}^{K}).$ (5)
 Hyperedge Weighting
@@
 member nodes) of each hyperedge, the Weighter will calculate each hyperedge’s weight, which is supposed to indicate the importance of that hyperedge, based on the member nodes and their features, i.e.
-$\bm{w}=\mathcal{F}_{w}(H)$.
+$\boldsymbol{w}=\mathcal{F}_{w}(H)$.
 These weights are obtained by a $\operatorname{Readout}$ module, which is composed of: 1) summing over all the non-masked nodes feature-wisely; 2) dim reduction operation.
-$w^{k}=\operatorname{Readout}(\bm{h}^{k})=\operatorname{DimReduction}({\bm{m}^{k}}^{T}\bm{h}^{k})\in\mathbb{R}$ (6)
+$w^{k}=\operatorname{Readout}(\boldsymbol{h}^{k})=\operatorname{DimReduction}({\boldsymbol{m}^{k}}^{T}\boldsymbol{h}^{k})\in\mathbb{R},$ (6)
 where $w^{k}$ is the weight of the $k$-th hyperedge and $\operatorname{DimReduction}$ is an MLP with ReLU activations, where the output dimension is $1$.
-For all hyperedges, we obtain $K$ hyperedges in total, $\bm{w}=[w^{1},w^{2},\cdots,w^{K}]^{T}\in\mathbb{R}^{K}$.
+For all hyperedges, we obtain $K$ hyperedges in total, $\boldsymbol{w}=[w^{1},w^{2},\cdots,w^{K}]^{T}\in\mathbb{R}^{K}$.
 Finally, these weights will be fed into the final linear head to predict the label of the instance,
-$\hat{Y}=\mathcal{F}_{l}(\bm{w})\in\mathbb{R}$ (7)
-In contrast to previous hypergraph construction methods (Jin et al., 2019; Xiao et al., 2019), which identify hyperedges by refining neighbors and assign weights by aggregating node features, HyBRiD makes these procedures learnable and thus is able to identify MIMR hyperedges in a data-driven way through expressive neural networks.
-We decide the number of hyperedges $K$ according to the study in Appendix E.2.
+$\hat{Y}=\mathcal{F}_{l}(\boldsymbol{w})\in\mathbb{R}.$ (7)
+In contrast to previous hypergraph construction methods (Jin et al., 2019; Xiao et al., 2019), which identify hyperedges by refining neighbors and simply aggregating node features, HyBRiD makes these procedures learnable and thus is able to identify MIMR hyperedges in a data-driven way through expressive neural networks.
+The number of hyperedges $K$ is decided according to the study in Appendix H.1.
+Computational Complexity
+The computational complexity of our model is $O(N^{2}K)$, which is just the same scale as that of MLPs even though we are addressing a more challenging task: identifying high-order relationships in an exponential space.
+Details of the complexity calculation can be found in Appendix C.
 4.2 Optimization Framework
-We propose a new IB framework named multi-head drop-bottleneck to optimize HyBRiD.
+Since there are no existing IB frameworks that can be applied in our context, we propose a new IB framework named multi-head drop-bottleneck to optimize HyBRiD.
 To adopt an information bottleneck view of HyBRiD, we consider $X$, $Y$ and $H$ are random variables in the Markovian chain $X\leftrightarrow Y\leftrightarrow H$.
 According to our MIMR objective, we optimize
-$\arg\max I(H;Y)-\beta I(H;X)$ (8)
-where $\beta$ is a coefficient trading off informativeness and redundancy.
+$\arg\max I(H;Y)-\beta I(H;X),$ (8)
+where $I(\cdot;\cdot)$ denotes the mutual information.
+$I(H;Y)$ corresponds to the informativeness and $I(H;X)$ corresponds to the redundancy.
+$\beta$ is a coefficient trading off informativeness and redundancy.
 Since optimizing the mutual information for high-dimensional continuous variables is intractable, we instead optimize the lower bound of Equation 8.
@@
 $=\mathbb{H}[Y]+\mathbb{E}_{p(Y,H)}[\log p(Y|H)]$
-$=\mathbb{H}[Y]+\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)]+\mathbb{E}_{p(H)}[\operatorname{KL}(p(Y|H)|q_{\phi}(Y|H))]$
-$\geq\mathbb{H}[Y]+\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)]$
+$=\mathbb{H}[Y]+\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)]$
+$\quad+\mathbb{E}_{p(H)}[\operatorname{KL}(p(Y|H)|q_{\phi}(Y|H))]$
+$\geq\mathbb{H}[Y]+\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)],$
 where $\mathbb{H}[\cdot]$ is the entropy computation.
@@
 $q_{\phi}$ can be considered as a model that predicts $Y$ based on $H$, which essentially corresponds to $\mathcal{F}_{l}\circ\mathcal{F}_{w}$, where $\circ$ is the function composition.
-In practice, we set $q_{\phi}$ as a Gaussian model with variance $1$ most probabilistic machine learning models for continuous data modeling.
-(Alemi et al., 2016; Luo et al., 2019; Peng et al., 2018; Kingma & Welling, 2013).
+In practice, we set $q_{\phi}$ as a Gaussian model with variance $1$ as most probabilistic machine learning models do for continuous data modeling (Alemi et al., 2016; Luo et al., 2019; Peng et al., 2018; Kingma & Welling, 2013).
 For the second term (redundancy) in Equation 8, we have
-Proposition 1.
+Proposition 4.1.
 (Upper bound of $I(H;X)$ in multi-head drop-bottleneck)
-$I(H;X)\leq\sum_{k=1}^{K}I(\bm{h}^{k};X)\leq\sum_{k=1}^{K}\sum_{i=1}^{N}I(\bm{h}^{k}_{i};X_{i})=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i})$ (10)
-where $\bm{h}^{i}_{j}$ and $X_{j}$ is the $j$-th row of $\bm{h}^{i}$ and $X$ respectively.
+$I(H;X)\leq\sum_{k=1}^{K}I(\boldsymbol{h}^{k};X)$ $\leq\sum_{k=1}^{K}\sum_{i=1}^{N}I(\boldsymbol{h}^{k}_{i};X_{i})$ (10)
+$=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i}),$
+where $\boldsymbol{h}^{k}_{i}$ and $X_{i}$ is the $i$-th row of $\boldsymbol{h}^{k}$ and $X$ respectively.
 $p^{k}_{\theta,i}$ is the mask probability in Equation 3.
@@
 The equality holds if and only if nodes are independent and hyperedges do not overlap.
-The second inequality is inspired by Kim et al.
-(2021).
+It’s important to clarify that optimizing Equation 10 does not imply a penalty for overlaps.
+The second inequality is inspired by (Kim et al., 2021).
 The proof of the proposition can be found in Appendix A.
-Therefore, instead of optimizing the intractable objective 8, we optimize its upper bound (i.e. loss function) according to Equation 9 and 10.
-$\mathcal{L}$ $=\|Y-\mathcal{F}_{l}\circ\mathcal{F}_{w}\circ\mathcal{F}_{c}(X)\|_{2}^{2}+\beta\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i})\geq-I(H,Y)+\beta I(H,X)$ (11)
-The learnable components are the shallow embeddings in Equation 3, the $\operatorname{DimReduction}$ MLP in Equation 6 and the LinearHead $\mathcal{F}_{l}$ in Equation 7.
-For how do we choose the trade-off coefficient $\beta$, see Appendix E.3 for more discussions.
-4.3 Computational Complexity
-Suppose we have $N$ regions as $N$ nodes and $K$ hyperedges.
-For the Constructor, the unignorable computation is from the mask operation.
-The computational complexity of this step for each hyperedge is $O(Nd)$ since it does a pairwise multiplication operation of a matrix of size $N\times d$.
-Given that there are $K$ hyperedges, the total complexity is $O(NKd)$.
-For the Weighter, the computation is from the dim reduction operation and the linear head.
-The dim reduction operation is an MLP.
-In this work, the hidden dimensions are a fraction of the original feature dimension.
-Therefore, the complexity of the dim-reduction MLP is $O(d^{2})$.
-The linear head only contributes $O(d)$, which is neglectable.
-As a result, the computational complexity of the whole model is $O(NKd+d^{2})=O(N^{2}K)$ since the feature dimension is equal to the number of regions (5.1).
-This complexity is just at the same scale as that of MLPs even though we are addressing a more challenging task: identifying high-order relationships in an exponential space.
+Therefore, instead of optimizing the intractable objective Equations 8, we optimize its upper bound (i.e. loss function) according to Equations 9 and 10.
+$\mathcal{L}$ $=\|Y-\mathcal{F}_{l}\circ\mathcal{F}_{w}\circ\mathcal{F}_{c}(X)\|_{2}^{2}+\beta\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i})$ (11)
+$\geq-I(H,Y)+\beta I(H,X).$
+In conclusion, The learnable components are the shallow embeddings in Equation 3, the $\operatorname{DimReduction}$ MLP in Equation 6 and the LinearHead $\mathcal{F}_{l}$ in Equation 7.
+For how we choose the trade-off coefficient $\beta$, see Appendix H.2 for more discussions.
 5 Experiments
-In this section, we conducted experiments to validate the quality of the learned hyperedges regarding the cognition phenotype outcome.
-Furthermore, we conducted ablation studies to validate the key components in our model.
-We also analyzed our results both quantitatively and qualitatively.
-5.1 Experiment Settings
+In this section, we conduct experiments to validate the quality of the learned hyperedges in terms of the predictive performance towards the cognition phenotype outcome.
+Furthermore, we conduct ablation studies to validate the key components in our model.
+We also analyze our results both quantitatively and qualitatively.
+5.1 Predictive Performance of Hypereges
+Datasets
+We consider two fMRI datasets:
+1) Autism Brain Imaging Data Exchange (ABIDE) (Craddock et al., 2013) is an open-source dataset.
+This dataset involved resting-state fMRI of patients from 17 international sites, as well as the anatomical and phenotypic data.
+Regions are obtained by Craddock 200 atlas (Craddock et al., 2012).
+We use the preprocessed version from the official website.
+For prediction targets, we choose three intelligence quotients: FIQ (full-scale intelligence), VIQ (verbal intelligence quotient), and PIQ (performance intelligence).
+2) Adolescent Brain Cognitive Development (ABCD) (Casey et al., 2018) is one of the largest public fMRI datasets.
+Access is limited and requires adherence to a rigorous data request procedure to acquire the data.
+The data is collected from $11,875$ children aged between 9 to 10 years old.
+The functional MRI (fMRI) data is collected from children when they were resting and when they performed three tasks (SST, EN-back, MID).
+We use the ABCD imaging data collected from the baseline (release 2.0) as well as the 2-year follow-up (release 3.0).
+In conclusion, we obtain $8$ sub-datasets (we refer to them as datasets from now on) from $2$ timepoints under $4$ tasks.
+Regions are obtained by AAL3v1 atlas (Rolls et al., 2020).
+For preprocess procedures of the ABCD dataset, please refer to Appendix D for more details.
+For the prediction target, we consider fluid intelligence as our label.
+Fluid intelligence reflects the general mental ability and plays a fundamental role in various cognitive functions.
+For region features, consistent with previous connectivity-based methods (Li et al., 2021; Kan et al., 2022b; Ktena et al., 2018; Said et al., 2023), we use a region’s Pearson correlation coefficients to all other regions as the region features.
+Other details of the data preprocessing and statistics of each dataset are summarized in Appendix D.
 Evaluation Metric
-To evaluate the quality of hyperedges obtained by HyBRiD, we use CPM (Shen et al., 2017), a standard model that could evaluate the correlation between connectivity and the prediction target, due to its high impact in the community.
+To evaluate the quality of hyperedges obtained by HyBRiD, we use CPM (Shen et al., 2017), a standard model that could evaluate the relevance between the connectivity and the prediction target, due to its high impact in the community.
 In the original implementation of CPM, weights of pairwise edges are obtained by Pearson correlation between nodes.
@@
 This process is formulated as
-$r^{\prime}=\operatorname{CPM}(\bm{w}_{p},Y)$ (12)
-where $\bm{w}_{p}\in\mathbb{R}^{K_{p}}$ denotes the pairwise edge weights and $K_{p}$ is the total number of pairwise edges.
-$r^{\prime}$ is a metric that measures the quality of weights based on positive and negative correlations to the cognition score.
-To evaluate the quality of the learned weights for our model and baselines, we concatenate the learned weights $\bm{w}_{h}\in\mathbb{R}^{K_{h}}$ with the pairwise ones together into the CPM, and thus adjust Equation 12 to
-$r=\operatorname{CPM}([\bm{w}_{p}\|\bm{w}_{h}],Y)$ (13)
+$r^{\prime}=\operatorname{CPM}(\boldsymbol{w}_{p},Y),$ (12)
+where $\boldsymbol{w}_{p}\in\mathbb{R}^{K_{p}}$ denotes the pairwise edge weights and $K_{p}$ is the total number of pairwise edges.
+$r^{\prime}$ is a metric that measures the quality of weights based on positive and negative correlations to the outcome.
+To evaluate the quality of the learned weights for our model and baselines, we concatenate the learned weights $\boldsymbol{w}_{h}\in\mathbb{R}^{K_{h}}$ with the pairwise ones together into the CPM, and thus adjust Equation 12 to
+$r=\operatorname{CPM}([\boldsymbol{w}_{p}\|\boldsymbol{w}_{h}],Y),$ (13)
 […]
-Datasets
-We use functional brain imaging data from the first release of the Adolescent Brain Cognitive Development (ABCD) study, collected from $11,875$ children aged between 9 to 10 years old (Casey et al., 2018).
-The functional MRI (fMRI) data is collected from children when they were resting and when they performed three emotional and cognitive tasks.
-The fMRI data is processed using BioImage Suite (Joshi et al., 2011) with the standard preprocessing procedures, such as slice time and motion correction, registration to the MNI template (see details in Greene et al.
-(2018) and Horien et al.
-(2019)).
-We delete scans with more than 0.10 mm mean frame-to-frame displacement.
-We use the ABCD imaging data collected from the baseline (release 2.0) as well as the 2-year follow-up (release 3.0).
-For each time point, we included children subjected to four conditions: the resting state where the brain is not engaged in any activity (Rest), the emotional n-back task (EN-back), the Stop Signal task (SST), and the Monetary Incentive Delay (MID) task.
-In conclusion, we obtain $8$ datasets from $2$ timepoints and $4$ conditions (Rest, SST, EN-back, MID).
-Statistics of each dataset are summarized in Appendix C.
-For the prediction target, we consider fluid intelligence as our label.
-Fluid intelligence reflects the general mental ability and plays a fundamental role in various cognitive functions.
-Recent literature has seen success in predicting intelligence based on pairwise predictive modeling (Dubois et al., 2018a; b), and we aim to improve the prediction accuracy of the current methods.
-Data preprocessing
-The raw fMRI data of an instance is represented in four dimensions (3 spatial dimensions + 1 temporal dimension), which can be imagined as a temporal sequence of 3D images.
-First, brain images are parceled into regions (or nodes) using the AAL3v1 atlas (Rolls et al., 2020).
-Following previous works (Kan et al., 2022b; Li et al., 2021; Thomas et al., 2022), each region’s time series is obtained by averaging all voxels in that region.
-Consistent with previous connectivity-based methods (Li et al., 2021; Kan et al., 2022b; Ktena et al., 2018; Said et al., 2023), for each region, we use its Pearson correlation coefficients to all regions as its features.
-We randomly split the data into train, validation, and test sets in a stratified fashion.
-The split ratio is 8:1:1.
-Baselines
-We compare our method with $3$ classes of baselines: 1) standard method, which is exactly the classical method that predicts outcomes based on pairwise edges (Shen et al., 2017; Dadi et al., 2019; Wang et al., 2021).
-The comparison with standard methods shows whether the high-order connectivity has its advantage over the classical pairwise one or not.
-2) hypergraph construction methods.
-We consider $k$NN (Huang et al., 2009), $l_{1}$ hypergraph (Wang et al., 2015), and $l_{2}$ hypergraph (Jin et al., 2019).
-3) Connectivity-based cognition prediction methods, which are state-of-the-art predictive models based on brain connectivity.
-We consider BrainNetGNN (Mahmood et al., 2021), BrainGNN (Li et al., 2021), and BrainNetTF (Kan et al., 2022b).
-BrainGB (Cui et al., 2022a) is a study of different brain graph neural network designs and we include its best design as a baseline.
-Note that since these models are unable to identify hyperedge structures of brain regions, we input their last layer embeddings (each entry as a weight) into the CPM model.
-Note that our weights $\bm{w}$ are also last layer embeddings in HyBRiD.
-Implementation & Training Details
-Due to the data scarcity, training on individual datasets would result in serious overfitting.
-To mitigate this, we train our model as well as baselines using all eight datasets together.
-Details of the overfitting issue and our parameter-sharing strategy are further discussed in Appendix E.1.
-See Appendix D for other implementation and training details such as hyperparameter choices and software/hardware specifications.
-5.2 Quality of Hyperedges
-We report $r$ values by CPM in Table 1.
-As we can see, HyBRiD outperforms the state-of-the-art predictive models on $7$ datasets, with an average improvement of $12.1\%$.
-Moreover, HyBRiD outperforms all other deep learning baselines in efficiency, with $87\%$ faster than the next fastest one (BrainNetTF).
-Refer to Appendix F for more runtime details.
+Type Model FIQ VIQ PIQ
+Standard pairwise $0.052$ $0.124$ $0.056$
+Hypergraph Construction $k$NN $0.023$ $0.093$ $0.056$
+$l_{1}$ hypergraph $0.043$ $0.125$ $0.061$
+$l_{2}$ hypergraph $0.148_{\pm 0.000}$ $0.141_{\pm 0.014}$ $0.063_{\pm 0.004}$
+Connectivity based Prediction BrainNetGNN $\underline{0.162_{\pm 0.042}}$ $\underline{0.199_{\pm 0.042}}$ $\underline{0.223_{\pm 0.025}}$
+BrainGB $0.125_{\pm 0.119}$ $0.154_{\pm 0.068}$ $0.157_{\pm 0.053}$
+BrainGNN $0.105_{\pm 0.041}$ $0.176_{\pm 0.049}$ $0.159_{\pm 0.051}$
+BrainNetTF $0.132_{\pm 0.111}$ $0.176_{\pm 0.053}$ $0.180_{\pm 0.054}$
+Ours HyBRiD $\mathbf{0.181_{\pm 0.040}}$ $\mathbf{0.204_{\pm 0.031}}$ $\mathbf{0.245_{\pm 0.042}}$
+Table 1: $r$ values of our hyperedges compared to baselines on the ABIDE dataset. Results are averaged over 10 runs. Deterministic methods do not have standard deviations.
 Type Model SST 1 EN-back 1 MID 1 Rest 1 SST 2 EN-back 2 MID 2 Rest 2
@@
 Ours HyBRiD $\mathbf{0.361_{\pm 0.058}}$ $\mathbf{0.348_{\pm 0.061}}$ $\mathbf{0.386_{\pm 0.060}}$ $\underline{0.223_{\pm 0.056}}$ $\mathbf{0.738_{\pm 0.054}}$ $\mathbf{0.714_{\pm 0.037}}$ $\mathbf{0.816_{\pm 0.053}}$ $\mathbf{0.730_{\pm 0.049}}$
-Table 1: $r$ values of our hyperedges compared to baselines. Results are averaged over 10 runs. Deterministic methods do not have standard deviations.
-Ablation Studies
-[…]
+Table 2: $r$ values of our hyperedges compared to baselines on the ABCD dataset. Results are averaged over 10 runs. Deterministic methods do not have standard deviations.
+Baselines
+We compare our method with $3$ classes of baselines: 1) standard method, which is exactly the classical method that predicts outcomes based on pairwise edges (Shen et al., 2017; Dadi et al., 2019; Wang et al., 2021).
+The comparison with standard methods shows whether the high-order connectivity has its advantage over the classical pairwise one or not.
+2) hypergraph construction methods.
+We consider $k$NN (Huang et al., 2009), $l_{1}$ hypergraph (Wang et al., 2015), and $l_{2}$ hypergraph (Jin et al., 2019).
+3) connectivity-based phenotypic prediction methods, which are state-of-the-art predictive models based on brain connectivity.
+We consider BrainNetGNN (Mahmood et al., 2021), BrainGNN (Li et al., 2021), and BrainNetTF (Kan et al., 2022b).
+BrainGB (Cui et al., 2022a) is a study of different brain graph neural network designs and we include its best design as a baseline.
+Since none of these models are able to identify hyperedge structures of brain regions, we input their last layer embeddings (each entry as a weight) into the CPM model.
+Note that our weights $\boldsymbol{w}$ are also last layer embeddings in HyBRiD.
 Model SST 1 EN-back 1 MID 1 Rest 1 SST 2 EN-back 2 MID 2 Rest 2
@@
 Table 3: Ablation studies on the masking mechanism. Results are averaged over 10 runs.
+Implementation & Training Details
+Hyperparameter choices and other details can be found in Appendix E.
+Results
+We report $r$ values by CPM in Table 1 and Table 2.
+As we can see, on the ABIDE dataset, HyBRiD consistently outperforms all the baselines on different targets, with an average improvement of $8.8\%$ compared to the state-of-the-art model.
+On the ABCD dataset, HyBRiD outperforms the state-of-the-art predictive models on $7$ datasets of ABCD, with an average improvement of $12.1\%$.
+The results demonstrate our model is able to learn informative hyperedges towards different phenotypic outcomes from fMRI data of various brain states.
+Rest 1 is the only dataset that our model reaches the second-best.
+We attribute this imperfection to the discrepancy between validation and test datasets.
+This is because, through the inspection of the training/validation set performances, we find our model performs better than BrainNetTF on the training and validation dataset.
+More discussions about it can be found in Appendix F.
+Runtime
+HyBRiD outperforms all other deep learning baselines in efficiency, with $87\%$ faster than the second-fastest one (BrainNetTF). Refer to Appendix I for more runtime details.
+Ablation Studies
+[…]
+Extra Experiments on Synthetic Dataset
+Since there are no ground-truth hyperedges in real-world datasets of learning informative hyperedges towards a specific outcome, we construct a synthetic dataset to verify if our model can recover the correct hyperedge structure under the MIMR objective.
+We use the precision, recall, and F1 score to measure the correctness of the learned hyperedges with respect to the ground truth.
+Although it is challenging to learn hyperedges when only supervised by the task label, we find that our model reaches high performances, with an average improvement of 28.3% in terms of the F1 score, compared to the strongest baselines.
+Details about the synthetic experiments can be found in Appendix G.
+Figure 4: Hyperedge profiles.
+(a) Hyperedge degree distribution of learned hyperedges.
+(b) Correlation between hyperedge degree and significance.
+(c) Comparison between the number of hyperedges and pairwise edges under different significance thresholds.
+The total number of hyperedges is $32$.
+And the total number of pairwise edges is $26,896$.
 5.2 Further Analysis
+In this subsection, we analyze the results of our model. We mainly use the ABCD dataset in the analysis since it is much larger than the ABIDE dataset.
 Hyperedge Degree Distribution
-We plot the hyperedge degree distribution in Fig.3(a).
+We plot the hyperedge degree distribution in Figure 3(a).
 We find there are two distinct clusters in the figure.
@@
 Hyperedges with Higher Degree are More Significant
-CPM conducts a significance test on pairwise edges and hyperedges internally based on a linear regression model, and thus we can obtain a P-value for each hyperedge from the significance test.
+Since CPM conducts a significance test on pairwise edges and hyperedges internally based on a linear regression model, we can obtain a P-value for each hyperedge from the significance test.
 We define the significance of a hyperedge as $1-P_{v}\in[0,1]$ where $P_{v}$ is the P-value of that hyperedge.
-Figure 4: Hyperedge profiles.
-(a) Hyperedge degree distribution of learned hyperedges.
-(b) Correlation between hyperedge degree and significance.
-(c) Comparison between the number hyperedges and pairwise edges under different significance thresholds.
-The total number of hyperedges is $32$.
-And the total number of pairwise edges is $26,896$.
-The relationship between hyperedge degree and its significance is shown in Fig 3(b).
+The relationship between hyperedge degree and its significance is shown in Figure  3(b).
 In this figure, we find a strong positive correlation between a hyperedge’s degree and its significance, which indicates that interactions of multiple brain regions play more important roles in cognition than pairwise or individual ones.
@@
 Region Importance
-Figure 5: Visualization of the frequency of each region under different conditions.
-To better understand the roles of each brain region in cognition under different conditions, we studied the frequency at which each region appears in a hyperedge out of all identified hyperedges.
+Figure 5: Visualization of the frequency of each region under different fMRI tasks.
+To better understand the roles of each brain region in cognition under different fMRI tasks, we study the frequency at which each region appears in a hyperedge out of all identified hyperedges.
 The frequency, which can be considered as a measure of region importance, is visualized in Figure 5.
-Visual regions (Fusiform, Cuneus, Calcarine) are especially active due to the intensive visual demands in all three conditions.
-We found that the Frontal_Inf_Oper and ACC_pre, recognized for their participation in response inhibition (Pornpattananangkul et al., 2016), frequently appear in the SST task.
+Visual regions (Fusiform, Cuneus, Calcarine) are especially active due to the intensive visual demands in all three tasks.
+We found that the Inferior frontal gyrus, opercular part and the Anterior cingulate cortex, pregenual, recognized for their participation in response inhibition (Pornpattananangkul et al., 2016), frequently appear in the SST task.
 This aligns with what the SST task was designed to test.
-Interestingly, of the three conditions (SST, EN-back, MID), only EN-back prominently involves the Amygdala, a key region for emotion processing.
-This makes sense as EN-back is the only condition related to emotion.
-More visualizations and interpretations on the resting-state region importance and individual hyperedges can be found in Appendix G.
+Interestingly, of the three tasks (SST, EN-back, MID), only EN-back prominently involves the Amygdala, a key region for emotion processing.
+This makes sense as EN-back is the only task related to emotion.
+More visualizations of individual hyperedges can be found in Appendix J.
 6 Conclusion
-In this work, we propose HyBRiD for identifying maximally informative yet minimally redundant (MIMR) high-order relationships of brain regions.
-To effectively optimize our model, we further proposed a novel information bottleneck framework with theoretical guarantees.
-Our method outperforms state-of-the-art models of hypergraph construction and connectivity-based prediction.
+In this work, we proposed HyBRiD for identifying maximally informative yet minimally redundant (MIMR) high-order relationships of brain regions.
+To effectively optimize our model, we further proposed a novel information bottleneck framework and derived its theory.
+Our method outperforms state-of-the-art models.
 The result analysis shows the effectiveness of our model.
-We expect such advancements could benefit clinical studies, providing insights into neurological disorders, and offering improved diagnostic tools in neurology.
+We expect such advancements could benefit clinical studies, providing insights into neurological disorders, and offering improved diagnostic tools in neurology and other related fields.
 Limitations
 HyBRiD only considers static high-order relations.
-Given that ABCD tasks are dynamic, including temporal changes and interactions, it will be interesting to study the evolution of these high-order relationships.
+Given that fMRI tasks are dynamic, including temporal changes and interactions, it will be interesting to study the evolution of these high-order relationships.
+7 Impact Statements
+This paper presents work whose goal is to advance the field of Machine Learning.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
 Appendix A Proof of the Upper Bound
@@
 Figure 6: The graphical model of random variables $X$, $Y$ and $Z$.
-Lemma 1.
+Lemma A.1.
 Given random variables $X$, $Y$ and $Z$. Their relationships are described in the graphical model illustrated in Figure 6. We have
@@
 Proof.
-$I(X;Y|Z)-I(X;Y)$ $=\int p(x,y,z)\log\frac{p(z)p(x,y,z)}{p(x,z)p(y,z)}\mathrm{d}x\mathrm{d}z\mathrm{d}y$ (15)
+$I(X;$ $Y|Z)-I(X;Y)$ (15)
+$=\int p(x,y,z)\log\frac{p(z)p(x,y,z)}{p(x,z)p(y,z)}\mathrm{d}x\mathrm{d}z\mathrm{d}y$
 $-\int p(x,y)\log\frac{p(x,y)}{p(x)p(y)}\mathrm{d}x\mathrm{d}y$
@@
 which finishes the proof.
-Corollary 1.1.
+Corollary A.2.
 Given the same graphical model 6, we have
@@
 $I(X,Z;Y)=I(X;Y|Z)+I(Z;Y)$ (17)
-According to Lemma 1, we have $I(X;Y|Z)\leq I(X;Y)$, which finishes the proof.
-Theorem 2.
+According to Lemma A.1, we have $I(X;Y|Z)\leq I(X;Y)$, which finishes the proof.
+Theorem A.3.
 For random variables in Equation 10, we have
-$I(H;X)\leq\sum_{k=1}^{K}I(\bm{h}^{k};X)$ (18)
+$I(H;X)\leq\sum_{k=1}^{K}I(\boldsymbol{h}^{k};X)$ (18)
 […]
-Figure 7: The graphical model of random variables $\bm{h}^{k}$ and $X$
+Figure 7: The graphical model of random variables $\boldsymbol{h}^{k}$ and $X$
 Proof.
 According to the definitions of $X$ and $H$, which are described in Section 4, we can draw a graphical model of them in Figure 7.
-Define a new random variable $\bm{h}^{k_{1}:k_{2}}=[\bm{h}^{k_{1}},\bm{h}^{k_{1}+1},\cdots,\bm{h}^{k_{2}-1},\bm{h}^{k_{2}}]$, which is a concatenation from $\bm{h}^{k_{1}}$ to $\bm{h}^{k_{2}}$.
-According to Corollary 1.1 we have
-$I(H;X)$ $\leq I(\bm{h}^{1},X)+I(\bm{h}^{2:K},X)$ (19)
-$\leq I(\bm{h}^{1},X)+I(\bm{h}^{2},X)+I(\bm{h}^{3:K},X)$
-$\leq I(\bm{h}^{1},X)+I(\bm{h}^{2},X)+I(\bm{h}^{3},X)+\cdots$
-$\leq\sum_{k=1}^{K}I(\bm{h}^{k};X)$
-Theorem 3.
+Define a new random variable $\boldsymbol{h}^{k_{1}:k_{2}}=[\boldsymbol{h}^{k_{1}},\boldsymbol{h}^{k_{1}+1},\cdots,\boldsymbol{h}^{k_{2}-1},\boldsymbol{h}^{k_{2}}]$, which is a concatenation from $\boldsymbol{h}^{k_{1}}$ to $\boldsymbol{h}^{k_{2}}$.
+According to Corollary A.2 we have
+$I(H;X)$ $\leq I(\boldsymbol{h}^{1},X)+I(\boldsymbol{h}^{2:K},X)$ (19)
+$\leq I(\boldsymbol{h}^{1},X)+I(\boldsymbol{h}^{2},X)+I(\boldsymbol{h}^{3:K},X)$
+$\leq I(\boldsymbol{h}^{1},X)+I(\boldsymbol{h}^{2},X)+I(\boldsymbol{h}^{3},X)+\cdots$
+$\leq\sum_{k=1}^{K}I(\boldsymbol{h}^{k};X)$
+Theorem A.4.
 (proposition 1)
-$I(H;X)\leq\sum_{k=1}^{K}\sum_{i=1}^{N}I(\bm{h}^{k}_{i};X_{i})=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i})$ (20)
+$I(H;X)\leq\sum_{k=1}^{K}\sum_{i=1}^{N}I(\boldsymbol{h}^{k}_{i};X_{i})=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i})$ (20)
 Proof.
 Given Theorem 2. It suffices to prove
-$I(\bm{h}^{k};X)\leq\sum_{i=1}^{N}I(\bm{h}^{k}_{i};X_{i})=\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i}),\quad\forall 1\leq k\leq K$ (21)
-And this is exactly the conclusion in Kim et al. (2021) if we consider $\bm{h}^{k}$ and $X$ as $X$ and $Z$ respectively in their paper. ∎
+$I(\boldsymbol{h}^{k};X)\leq\sum_{i=1}^{N}I(\boldsymbol{h}^{k}_{i};X_{i})=\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i}),\quad\forall 1\leq k\leq K$ (21)
+And this is exactly the conclusion in (Kim et al., 2021) if we consider $\boldsymbol{h}^{k}$ and $X$ as $X$ and $Z$ respectively in their paper. ∎
 Appendix B Connectome-Based Predictive Modeling
@@
 (2017); Finn et al.
-(2015)) has shown tremendous promise in recent years in detecting imaging biomarkers.
-(Rosenberg et al., 2015; Dubois et al., 2018a; Rosenberg et al., 2020; 2016).
+(2015) have shown tremendous promise in recent years in detecting imaging biomarkers by CPM (connectome-based predictive modeling) (Rosenberg et al., 2015; Dubois et al., 2018; Rosenberg et al., 2020, 2016).
 Such a model, based chiefly on functional MRI data, can measure the significance of the input edge weights, which is revealed by a correlation coefficient that reflects the correlation between the edge weights and the neurological outcomes.
@@
 Note that this correlation coefficient is equivalent to the one between the summed weights and the outcomes, and is exactly the metric $r$ we use to evaluate our hyperedges in Equation 13.
-Since positive edges and negative edges will cancel out with each other when being summed, we adopt the combining strategy in Boyle et al. (2023).
+Since positive edges and negative edges will cancel out with each other when being summed, we adopt the combining strategy in (Boyle et al., 2023).
 CPM Measures the Quality of Edge Weights
-According to step 6, the evaluation of the predictive model could be measured by the correlation between predicted and ground-truth outcomes Shen et al.
-(2017).
-Since CPM is a linear model that predicts the outcome based on the sum of significant edge weights, the correlation is equal to the correlation between the sum and ground-truth outcomes (which is exactly the r in Eq.
-13).
+According to step 6, the evaluation of the predictive model could be measured by the correlation between predicted and ground-truth outcomes (Shen et al., 2017).
+Since CPM is a linear model that predicts the outcome based on the sum of significant edge weights, the correlation is equal to the correlation between the sum and ground-truth outcomes (which is exactly the $r$ in Equation 13).
 Hence, one can expect a larger correlation if the edge weights are more correlated (and thus are more predictive).
@@
 where $n$ is the number of samples and $B$ is the beta function. Based on the distribution, we obtain a $P$-value for the $k$-th edge, which is used to measure the significance of the edge.
-Appendix C Dataset details
+Appendix C Computational Complexity
+Suppose we have $N$ regions as $N$ nodes and $K$ hyperedges.
+For the Constructor, the unignorable computation is from the mask operation.
+The computational complexity of this step for each hyperedge is $O(Nd)$ since it does a pairwise multiplication operation of a matrix of size $N\times d$.
+Given that there are $K$ hyperedges, the total complexity is $O(NKd)$.
+For the Weighter, the computation is from the dim reduction operation and the linear head.
+The dim reduction operation is an MLP.
+In this work, the hidden dimensions are a fraction of the original feature dimension.
+Therefore, the complexity of the dim-reduction MLP is $O(d^{2})$.
+The linear head only contributes $O(d)$, which is neglectable.
+As a result, the computational complexity of the whole model is $O(NKd+d^{2})=O(N^{2}K)$ since the feature dimension is equal to the number of regions (See Appendix D for features we used).
+This complexity is just at the same scale as that of MLPs even though we are addressing a more challenging task: identifying high-order relationships in an exponential space.
+Appendix D Dataset details
+D.1 Preprocessing of ABCD Dataset
+We use the preprocessed ABIDE dataset from the official website. We preprocessed the restricted ABCD dataset ourselves. Below is the preprocessing procedure of the ABCD dataset.
+raw data to voxel-level fMRI time series
+The fMRI data is processed using BioImage Suite (Joshi et al., 2011).
+First, we performed motion correction and slice-time correction using SPM5; and via BioImage Suite, the data were registered to a standardized $3mm\times 3mm\times 3mm$ common space, where we generated masks representing white matter, gray matter, and cerebrospinal fluid (CSF) and computed the mean time courses for both white matter and CSF.
+We orthogonalized each gray matter time course with respect to the mean time courses of both white matter and CSF, and we orthogonalized each gray matter time course to the six motion-related signals via SPM5.
+We then applied a bandpass Butterworth filter with a frequency range of $0.02$Hz to $0.1$Hz to the orthogonalized time courses.
+We used a Gaussian kernel with a full-width at half-maximum (FWHM) of $6$mm to enhance spatial coherence and spatial smoothing.
+Lastly, we removed the linear trend from all signals in accordance with the methodology detailed in (Shen et al., 2013).
+We deleted scans with more than 0.10 mm mean frame-to-frame displacement.
+Additional details about the standard preprocessing procedures, such as slice time and motion correction, and registration to the MNI template can be found in (Greene et al., 2018) and (Horien et al., 2019).
+voxel-level fMRI time series to region features
+The fMRI time series data of a human subject is represented in four dimensions (3 spatial dimensions + 1 temporal dimension), which can be imagined as a temporal sequence of 3D images.
+First, brain images are parceled into regions (or nodes) using the AAL3v1 atlas (Rolls et al., 2020).
+Following previous works (Kan et al., 2022b; Li et al., 2021; Thomas et al., 2022), each region’s time series is obtained by averaging all voxels in that region.
+Consistent with previous connectivity-based methods (Li et al., 2021; Kan et al., 2022b; Ktena et al., 2018; Said et al., 2023), for each region, we use its Pearson correlation coefficients to all regions as its features.
+We randomly split the data into train, validation, and test sets in a stratified fashion.
+The split ratio is 8:1:1.
+D.2 Dataset Statistics
 The statistics of the number of instances and the time series length are summarized in Table 4.
-Dataset Rest 1 SST 1 EN-back 1 MID 1 Rest 2 SST 2 EN-back 2 MID 2
-#(instances) 1676 1673 1678 1678 1949 1053 1044 1062
-length of time series 375 437 362 403 375 437 362 403
-Table 3: Statistics of 8 datasets.
-Appendix D Training Details
+Dataset ABIDE ABCD
+FIQ VIQ PIQ Rest 1 SST 1 EN-back 1 MID 1 Rest 2 SST 2 EN-back 2 MID 2
+#(instances) 1035 1035 1035 1676 1673 1678 1678 1949 1053 1044 1062
+length of time series 196 196 196 375 437 362 403 375 437 362 403
+Table 4: Statistics of datasets we use.
+ABIDE dataset only contains resting-state fMRI data.
+ABCD contains resting-state and task-based data.
+We use ABCD of $2$ timepoints.
+For example, Rest 1 means resting-state fMRI data from timepoint $1$.
+Appendix E Training Details
+Due to the data scarcity, we found training on single sub-datasets of the ABCD dataset leads to severe overfitting.
+To mitigate this, we train our model as well as all the baselines on all datasets together and report the results individually.
+Note that on the ABIDE dataset, we train our model under the three targets separately since we don’t encounter such an issue on the ABIDE dataset.
 Hardware
-We train our model on a machine with a Intel Xeon Gold 6326 CPU and RTX A5000 GPUs.
+We train our model on a machine with an Intel Xeon Gold 6326 CPU and RTX A5000 GPUs.
 Software
-See Table 4 for software we used and the versions.
+See Table 5 for the software we used and the versions.
 software version
@@
 nibabel 4.0.2
-Table 4: Software versions
+Table 5: Software versions.
 Hyperparameter Choices
-The hyperparameters selection is shown in Table 5. Some crucial hyperparameters ablation experiments can be found in Appendix E.
-notation meaning value
-$lr$ learning rate $1\times 10^{-3}$
-$K$ number of hyperedges $32$
-$\beta$ trade-off coefficients information bottleneck $0.2$
-$[h_{1},h_{2},h_{3}]$ hidden sizes of the dim reduction MLP $[32,8,1]$
-$B$ batch size $64$
-Table 5: Hyperparameter choices
-Appendix E More Ablation Studies
-E.1 Parameters Sharing
-As previously discussed, the scarcity of fMRI data can result in serious over-fitting.
-In our experiments, we observed that when the model is trained on each dataset individually, the validation set loss remains stagnant, which indicates over-fitting.
-To address this, we examined the model’s performance under various parameter-sharing strategies.
-Specifically, we have three learnable components: the shallow embeddings in Equation 3, the dim reduction MLP in 6 and the LinearHead $\mathcal{F}_{l}$.
-These components can be either shared or condition-specific.
-Results of different parameter-sharing strategy is shown in Table 6.
-Strategy ID ShallowEmbeddings DimReduction LinearHead Average Score
-1 $0.337$
-2 $0.335$
-3 $0.508$
-4 $\underline{0.539}$
-5 $0.351$
-6 $0.384$
-7 $0.506$
-8 $\mathbf{0.558}$
-Table 6: The table summarizes the model’s average performance under various parameter-sharing strategies.
-The symbol ✓ indicates a component being condition-specific, while an empty cell implies a component being shared across conditions.
-Note that strategy 1 is equivalent to training models separately for each condition.
-From the table, we find that the DimReduction component plays the most significant role.
-Making DimReduction condition-specific enhances the average performance by $10\%-20\%$.
-Strategy 8 is the best.
-However, domain experts may expect different structures under different conditions.
-Therefore, we still adopt condition-specific ShallowEmbeddings (i.e.
-strategy 4), given that their scores are really close.
-We argue that the fact that strategy 4 is not better than strategy 8 is because our prediction target is fluid intelligence, which is a general measurement of cognition and thus tends to be condition-agnostic.
-We anticipate a potential shift in results if we switch to a more condition-specific cognition phenotype outcome.
-E.2 Choices of number of hyperedges $K$
+The hyperparameters selection is shown in Table 6. Some crucial hyperparameters ablation experiments can be found in Appendix H.
+notation
+meaning
+value
+$lr$
+learning rate
+$1\times 10^{-3}$
+$K$
+number of hyperedges
+$32$
+$\beta$
+trade-off coefficients information bottleneck
+$0.2$
+$[h_{1},h_{2},h_{3}]$
+hidden sizes of the dim reduction MLP
+$[32,8,1]$
+$B$
+batch size
+$64$
+Table 6: Hyperparameter choices.
+Appendix F Why is HyBRiD not the Best on the Rest 1 Dataset?
+Our model reaches the second-best on the Rest 1 dataset.
+To find out the reason, we inspect the performance on both the training, validation and test datasets.
+We summarize the performances of our model and the SOTA one in the table below
+model train validation test
+BrainNetTF $0.967$ $0.484$ $0.334$
+HyBRiD (Ours) $0.984$ $0.491$ $0.223$
+Table 7: Our model’s performance compared to the SOTA one on the Rest 1 dataset.
+We find our model outperforms BrainNetTF on both training and validation datasets but fails on test datasets.
+Therefore, given that Rest 1 is the most noisy dataset, we argue that the imperfection is because
+1) Our model overfits the training dataset more than BrainNetTF.
+2) The validation score is higher, indicating the reason is the discrepancy between the training dataset and the test dataset.
+Appendix G Evaluation on Synthetic Dataset
+Dataset Synthesis
+Metric $K=1$ $K=5$ $K=10$ $K=30$
+P R F1 P R F1 P R F1 P R F1
+Random $0.158$ $0.094$ $0.110$ $0.182$ $0.170$ $0.171$ $0.224$ $0.218$ $0.218$ $\underline{0.288}$ $0.250$ $0.266$
+$k$NN $0.533$ $\mathbf{1.000}$ $0.696$ $0.247$ $0.241$ $0.243$ $0.144$ $0.156$ $0.148$ $0.139$ $0.128$ $0.133$
+$l_{1}$ hypergraph $0.160$ $\mathbf{1.000}$ $0.276$ $0.180$ $0.016$ $0.028$ $0.135$ $0.016$ $0.028$ $0.256$ $0.026$ $0.047$
+$l_{2}$ hypergraph $\underline{0.987}$ $0.825$ $\underline{0.897}$ $\underline{0.412}$ $\underline{0.631}$ $\underline{0.494}$ $\underline{0.276}$ $\underline{0.499}$ $\underline{0.351}$ $0.233$ $\underline{0.470}$ $0.310$
+HGCN $0.625$ $\mathbf{1.000}$ $0.769$ $0.196$ $0.477$ $0.278$ $0.188$ $0.379$ $0.251$ $0.277$ $0.364$ $\underline{0.314}$
+HyBRiD (Ours) $\mathbf{1.000}$ $\mathbf{1.000}$ $\mathbf{1.000}$ $\mathbf{0.595}$ $\mathbf{0.680}$ $\mathbf{0.631}$ $\mathbf{0.504}$ $\mathbf{0.604}$ $\mathbf{0.549}$ $\mathbf{0.428}$ $\mathbf{0.393}$ $\mathbf{0.401}$
+Table 8: Precision, recall and F1 score on synthetic dataset under different $K$, where $K$ is the number of hyperedges.
+The performance of the Random method increases as $K$ increases because the Hungarian algorithm is likely to provide better matches when there are more candidates.
+Our synthetic dataset is constructed as follows:
+1) Structure Generation: For each hyperedge, we randomly sample the hyperedge degree $d$ from a discrete uniform distribution between $2$ and $d_{\text{max}}$.
+After that, we randomly sample $d$ nodes as members of this hyperedge.
+2) Feature generation: For each hyperedge, we randomly sample a scalar value $v$ from the uniform distribution $U(0,1)$.
+We then randomly sample the features from $U(0,2v)$ for each node in this hyperedge.
+3) Label Generation: For each hyperedge, we calculate the maximum of its node features as the summary of the hyperedge.
+We sum the summaries of all hyperedges to get a single value $Y$ as the label of the hypergraph.
+Note that consistent with our settings of real-world datasets, the structure is shared across all hypergraphs, while node features and hyperedge weights are different.
+Metrics
+We use the macro-precision, recall and F1 score to measure the correctness of the learned hyperedges with respect to the ground-truth ones.
+Note that when learning on the synthetic dataset, the order of hyperedge may differ from the ground truths.
+Therefore, the Hungarian algorithm is employed to match the learned hyperedges with the ground truth.
+The precision, recall and F1 score are calculated after matching.
+Baselines
+We use the hyperedge construction methods in Section 5, i.e.
+$k$NN(Huang et al., 2009), $l_{1}$ hypergraph(Wang et al., 2015) and $l_{2}$ hypergraph(Jin et al., 2019) as our baselines.
+Besides, we implement a hypergraph structure learning model.
+This is a hypergraph convolutional neural network (HGCN) (Bai et al., 2021), where both the parameters of layers and the hypergraph structure (the incidence matrix) are learnable.
+Implementaton Details
+The maximum degree $d_{\text{max}}$ is set to $34$, which is the maximum degree of hyperedges of the ABCD dataset according to the analysis in Section 5.
+The number of nodes is set to $164$, which is the number of regions of AAL3v1 atlas, used in the ABCD experiment.
+We conduct the experiments under different numbers of hyperedges (i.e.
+$K=1,5,10,30$).
+Results
+We report the results under different numbers of hyperedges.
+Although this is a hard task, our model consistently outperforms all the baselines significantly, with an average improvement of 28.3% in terms of the F1 score.
+This demonstrates that our model can learn the hyperedges well under the MIMR objective, without the direct supervision of hyperedge ground truths.
+Appendix H More Ablation Studies
+H.1 Choices of Number of Hyperedges $K$
 As explained in Section 4, we use $K$ heads for $K$ hyperedges. We study the correlation between the $r$ value and the number of hyperedges on three datasets:
@@
 (b) Performance with increasing $\beta$ from $0.1$ to $1.0$ with step $0.1$ on all datasets.
+Figure 8: Studies on the choice of two key hyperparameters, $\beta$ and $K$, in our model.
 From Figure 7(a), we find that the overall performance increases dramatically before $K=17$, but becomes stable and close to saturation after $K=32$.
 To improve the efficiency while ensuring the performance, we choose $K=32$.
-E.3 Choices of the trade-off coefficient $\beta$
+H.2 Choices of the Trade-off Coefficient $\beta$
 In our optimization objective 8, $\beta$ acts as a trade-off parameter, which is a non-negative scalar that determines the weight given to the second term relative to the first.
@@
 Accordingly, we adopt $\beta=0.2$.
-Appendix F Runtime Comparison
+Appendix I Runtime Comparison
 Figure 9 summarizes the per-batch training time of all deep learning models.
@@
 Figure 9: Training time per batch of all deep learning models.
-Appendix G More Visualizations
+Figure 10: Visualization of the most significant hyperedge of the EN-back task.
+Appendix J More Visualizations
+Resting-State Brain Region Importance
+We visualize the region importance of the resting state in Figure 11.
+Different from task states, where specific brain regions are activated in response to particular tasks, brain activities during resting states, are not driven by external tasks, leading to more diffuse and less predictable patterns of activation.
+This makes it harder to pinpoint specific interactions or functions.
+Figure 11: Region importance of resting state.
 Hyperedge Case Study
-We visualize the most significant hyperedge under the EN-back condition.
+We visualize the most significant hyperedge of the EN-back task in Figure 10.
 We observe a coordinated interaction of numerous brain regions, each fulfilling specific roles.
@@
   • Visual Processing: Calcarine_R, Lingual_L, Fusiform_L.
-These regions are responsible for visual perception and some of them are related to complex visual contents like symbols and human faces, which were presented during the task.
+These regions are responsible for visual perception and some of them are related to complex visual contents like symbols and human faces, which were presented during the fMRI task.
   • Sensory SupraMarginal_L: It is responsible for interpreting tactile sensors and perceiving limbs location.
@@
   • Cognitive Control ACC_pre_L: In the EN-back task, this region is likely crucial for maintaining focus, error detection and correction, conflict management in working memory, and modulating emotional responses to the task’s demands.
-Figure 9: Visualization of the most significant hyperedge in EN-back condition
-Resting-state brain region importance
-We visualize the region importance of the resting state in Figure 10.
-Different from task states, where specific brain regions are activated in response to particular tasks, brain activities during resting states, are not driven by external tasks, leading to more diffuse and less predictable patterns of activation.
-This makes it harder to pinpoint specific interactions or functions.
-Figure 10: Region importance of resting state.
-body>

### v3

@@
 Our model outperforms the state-of-the-art predictive model by an average of $11.2\%$, regarding the quality of hyperedges measured by CPM, a standard protocol for studying brain connections.
+Source code is available at https://github.com/Graph-and-Geometric-Learning/HyBRiD.
 Machine Learning, ICML
@@
 For instance, executive planning requires the appropriate communication of signals across many distinct cortical areas (Logue & Gould, 2014).
-Pairwise connectivity, without considering the brain’s high-order structure, can lead to inconsistent findings across studies with low prediction performance.
+These high-order relationships, cannot always be decomposed into pairwise ones (Battiston et al., 2020, 2021; Bick et al., 2023), thus allowing them to capture information beyond the reach of pairwise ones (Do et al., 2020).
+Consequently, relying solely on pairwise connectivity, without accounting for the brain’s complex high-order structure, may result in inconsistent findings and low predictive performance across studies.
 Although recently there have been a few works (Zu et al., 2016; Xiao et al., 2019; Li et al., 2022) working on discovering the high-order relationships of brain regions, they are unable to effectively extract meaningful patterns.
@@
 As a result, most informative high-order relations are likely excluded at the first stage.
-This inspires us to solve the problem in an end-to-end manner through a more expressive model.
+This inspires us to solve the problem in an end-to-end manner through a more expressive model and a more appropriate objective.
 In this paper, we aim to identify high-order relationships that are informative towards a phenotypic outcome, such as a cognition score.
@@
 However, current methods for hypergraph construction, which are mostly based on neighbor distances and neighbor reconstruction (Wang et al., 2015; Liu et al., 2017; Jin et al., 2019; Huang et al., 2009), are unsuitable in our context for several reasons: 1) they are unable to learn MIMR hyperedges due to the absence of a tractable objective for learning such hyperedges.
-2) they fall short of learning consistent structures across subjects, which contradicts the belief that the cognitive function mechanism of healthy human beings should be similar (Wang et al., 2023).
+2) they fall short of learning consistent structures across subjects, which contradicts the belief that the cognitive function mechanism of healthy human beings should be similar (Wang et al., 2023b).
 3) the number of hyperedges is restricted to the number of nodes, which may lead to sub-optimal performance.
@@
 However, the number of hyperedges of these methods is restricted to the number of nodes, and hyperedges obtained by these methods are inconsistent across instances.
+Cluster-based approaches learn community structure in a graph (Bannadabhavi et al., 2023; Ying et al., 2018), which can also be used to form hyperedges.
 Zhang et al.
@@
 Information Bottleneck
-Information bottleneck (IB) (Tishby et al., 2000) is originally a technique in data compression.
+Information bottleneck (IB) (Tishby et al., 2000) is a technique in data compression.
 The key idea is to extract a summary of data, which contains the most relevant information to the objective.
@@
 In contrast to previous hypergraph construction methods (Jin et al., 2019; Xiao et al., 2019), which identify hyperedges by refining neighbors and simply aggregating node features, HyBRiD makes these procedures learnable and thus is able to identify MIMR hyperedges in a data-driven way through expressive neural networks.
-The number of hyperedges $K$ is decided according to the study in Appendix H.1.
+The number of hyperedges $K$ is decided according to the study in Appendix I.1.
 Computational Complexity
@@
 In conclusion, The learnable components are the shallow embeddings in Equation 3, the $\operatorname{DimReduction}$ MLP in Equation 6 and the LinearHead $\mathcal{F}_{l}$ in Equation 7.
-For how we choose the trade-off coefficient $\beta$, see Appendix H.2 for more discussions.
+For how we choose the trade-off coefficient $\beta$, see Appendix I.2 for more discussions.
 5 Experiments
@@
 $r^{\prime}$ is a metric that measures the quality of weights based on positive and negative correlations to the outcome.
-To evaluate the quality of the learned weights for our model and baselines, we concatenate the learned weights $\boldsymbol{w}_{h}\in\mathbb{R}^{K_{h}}$ with the pairwise ones together into the CPM, and thus adjust Equation 12 to
-$r=\operatorname{CPM}([\boldsymbol{w}_{p}\|\boldsymbol{w}_{h}],Y),$ (13)
-[…]
+To evaluate the quality of the learned weights for our model, we replace the pairwise edge weights with the learned high-order weights $\boldsymbol{w}_{h}\in\mathbb{R}^{K_{h}}$, and thus adjust Equation 12 to
+$r=\operatorname{CPM}(\boldsymbol{w}_{h},Y),$ (13)
+Comparing $r$ to $r^{\prime}$ reflects the quality of learned weights in terms of the prediction performance since it measures the overall correlation between weights and the prediction target.
+In our model, $\boldsymbol{w}_{h}=\boldsymbol{w}$, which is the learned hyperedge weights.
+Table 1: $r$ values of our hyperedges compared to baselines on the ABCD dataset. Results are averaged over 10 runs. Deterministic methods do not have standard deviations.
+Type Model SST 1 EN-back 1 MID 1 Rest 1 SST 2 EN-back 2 MID 2 Rest 2
+Standard pairwise $0.113$ $0.218$ $0.099$ $0.164$ $0.201$ $0.322$ $0.299$ $0.289$
+Hypergraph Construction $k$NN $0.115$ $0.268$ $0.168$ $0.127$ $0.257$ $0.266$ $0.238$ $0.315$
+$l_{1}$ hypergraph $0.099$ $0.223$ $0.125$ $0.126$ $0.145$ $0.295$ $0.242$ $0.259$
+$l_{2}$ hypergraph $0.096_{\pm 0.002}$ $0.197_{\pm 0.003}$ $0.118_{\pm 0.003}$ $0.157_{\pm 0.016}$ $0.203_{\pm 0.005}$ $0.272_{\pm 0.004}$ $0.289_{\pm 0.011}$ $0.307_{\pm 0.006}$
+Connectivity based Prediction BrainNetGNN $0.227_{\pm 0.060}$ $0.287_{\pm 0.043}$ $0.266_{\pm 0.046}$ $0.221_{\pm 0.040}$ $0.468_{\pm 0.058}$ $0.480_{\pm 0.068}$ $0.506_{\pm 0.057}$ $0.453_{\pm 0.028}$
+BrainGB $0.190_{\pm 0.073}$ $0.214_{\pm 0.051}$ $0.265_{\pm 0.048}$ $0.176_{\pm 0.066}$ $0.447_{\pm 0.089}$ $0.483_{\pm 0.077}$ $0.458_{\pm 0.064}$ $0.432_{\pm 0.076}$
+BrainGNN $0.262_{\pm 0.030}$ $0.235_{\pm 0.032}$ $0.260_{\pm 0.049}$ $0.185_{\pm 0.058}$ $0.455_{\pm 0.028}$ $0.391_{\pm 0.077}$ $0.445_{\pm 0.078}$ $0.368_{\pm 0.041}$
+BrainNetTF $\underline{0.327_{\pm 0.084}}$ $\underline{0.338_{\pm 0.056}}$ $\underline{0.370_{\pm 0.098}}$ $\mathbf{0.334_{\pm 0.084}}$ $\underline{0.633_{\pm 0.178}}$ $\underline{0.631_{\pm 0.142}}$ $\underline{0.629_{\pm 0.123}}$ $\underline{0.588_{\pm 0.138}}$
+Ours HyBRiD $\mathbf{0.361_{\pm 0.058}}$ $\mathbf{0.348_{\pm 0.061}}$ $\mathbf{0.386_{\pm 0.060}}$ $\underline{0.223_{\pm 0.056}}$ $\mathbf{0.738_{\pm 0.054}}$ $\mathbf{0.714_{\pm 0.037}}$ $\mathbf{0.816_{\pm 0.053}}$ $\mathbf{0.730_{\pm 0.049}}$
+Table 2: $r$ values of our hyperedges compared to baselines on the ABIDE dataset. Results are averaged over 10 runs. Deterministic methods do not have standard deviations.
 Type Model FIQ VIQ PIQ
@@
 Ours HyBRiD $\mathbf{0.181_{\pm 0.040}}$ $\mathbf{0.204_{\pm 0.031}}$ $\mathbf{0.245_{\pm 0.042}}$
-Table 1: $r$ values of our hyperedges compared to baselines on the ABIDE dataset. Results are averaged over 10 runs. Deterministic methods do not have standard deviations.
-Type Model SST 1 EN-back 1 MID 1 Rest 1 SST 2 EN-back 2 MID 2 Rest 2
-Standard pairwise $0.113$ $0.218$ $0.099$ $0.164$ $0.201$ $0.322$ $0.299$ $0.289$
-Hypergraph Construction $k$NN $0.115$ $0.268$ $0.168$ $0.127$ $0.257$ $0.266$ $0.238$ $0.315$
-$l_{1}$ hypergraph $0.099$ $0.223$ $0.125$ $0.126$ $0.145$ $0.295$ $0.242$ $0.259$
-$l_{2}$ hypergraph $0.096_{\pm 0.002}$ $0.197_{\pm 0.003}$ $0.118_{\pm 0.003}$ $0.157_{\pm 0.016}$ $0.203_{\pm 0.005}$ $0.272_{\pm 0.004}$ $0.289_{\pm 0.011}$ $0.307_{\pm 0.006}$
-Connectivity based Prediction BrainNetGNN $0.227_{\pm 0.060}$ $0.287_{\pm 0.043}$ $0.266_{\pm 0.046}$ $0.221_{\pm 0.040}$ $0.468_{\pm 0.058}$ $0.480_{\pm 0.068}$ $0.506_{\pm 0.057}$ $0.453_{\pm 0.028}$
-BrainGB $0.190_{\pm 0.073}$ $0.214_{\pm 0.051}$ $0.265_{\pm 0.048}$ $0.176_{\pm 0.066}$ $0.447_{\pm 0.089}$ $0.483_{\pm 0.077}$ $0.458_{\pm 0.064}$ $0.432_{\pm 0.076}$
-BrainGNN $0.262_{\pm 0.030}$ $0.235_{\pm 0.032}$ $0.260_{\pm 0.049}$ $0.185_{\pm 0.058}$ $0.455_{\pm 0.028}$ $0.391_{\pm 0.077}$ $0.445_{\pm 0.078}$ $0.368_{\pm 0.041}$
-BrainNetTF $\underline{0.327_{\pm 0.084}}$ $\underline{0.338_{\pm 0.056}}$ $\underline{0.370_{\pm 0.098}}$ $\mathbf{0.334_{\pm 0.084}}$ $\underline{0.633_{\pm 0.178}}$ $\underline{0.631_{\pm 0.142}}$ $\underline{0.629_{\pm 0.123}}$ $\underline{0.588_{\pm 0.138}}$
-Ours HyBRiD $\mathbf{0.361_{\pm 0.058}}$ $\mathbf{0.348_{\pm 0.061}}$ $\mathbf{0.386_{\pm 0.060}}$ $\underline{0.223_{\pm 0.056}}$ $\mathbf{0.738_{\pm 0.054}}$ $\mathbf{0.714_{\pm 0.037}}$ $\mathbf{0.816_{\pm 0.053}}$ $\mathbf{0.730_{\pm 0.049}}$
-Table 2: $r$ values of our hyperedges compared to baselines on the ABCD dataset. Results are averaged over 10 runs. Deterministic methods do not have standard deviations.
+Table 3: Ablation studies on the masking mechanism. Results are averaged over 10 runs.
+Model SST 1 EN-back 1 MID 1 Rest 1 SST 2 EN-back 2 MID 2 Rest 2
+HyBRiD $\mathbf{0.361_{\pm 0.058}}$ $\mathbf{0.348_{\pm 0.061}}$ $\mathbf{0.386_{\pm 0.060}}$ $\underline{0.223_{\pm 0.056}}$ $\mathbf{0.738_{\pm 0.054}}$ $\mathbf{0.714_{\pm 0.037}}$ $\mathbf{0.816_{\pm 0.053}}$ $\mathbf{0.730_{\pm 0.049}}$
+HyBRiDNoMask $0.297_{\pm 0.035}$ $0.274_{\pm 0.057}$ $\underline{0.323_{\pm 0.059}}$ $0.221_{\pm 0.034}$ $0.653_{\pm 0.036}$ $0.599_{\pm 0.059}$ $0.757_{\pm 0.021}$ $0.543_{\pm 0.038}$
+HyBRiDRndMask $0.256_{\pm 0.069}$ $0.191_{\pm 0.046}$ $0.255_{\pm 0.080}$ $0.190_{\pm 0.051}$ $0.541_{\pm 0.069}$ $0.514_{\pm 0.038}$ $0.598_{\pm 0.064}$ $0.482_{\pm 0.083}$
+HyBRiDSoftMask $\underline{0.343_{\pm 0.042}}$ $\underline{0.314_{\pm 0.040}}$ $0.320_{\pm 0.055}$ $\mathbf{0.245_{\pm 0.061}}$ $\underline{0.707_{\pm 0.042}}$ $\underline{0.662_{\pm 0.058}}$ $\underline{0.796_{\pm 0.031}}$ $\underline{0.655_{\pm 0.030}}$
 Baselines
@@
 Note that our weights $\boldsymbol{w}$ are also last layer embeddings in HyBRiD.
-Model SST 1 EN-back 1 MID 1 Rest 1 SST 2 EN-back 2 MID 2 Rest 2
-HyBRiD $\mathbf{0.361_{\pm 0.058}}$ $\mathbf{0.348_{\pm 0.061}}$ $\mathbf{0.386_{\pm 0.060}}$ $\underline{0.223_{\pm 0.056}}$ $\mathbf{0.738_{\pm 0.054}}$ $\mathbf{0.714_{\pm 0.037}}$ $\mathbf{0.816_{\pm 0.053}}$ $\mathbf{0.730_{\pm 0.049}}$
-[…]
-[…]
-[…]
-Table 3: Ablation studies on the masking mechanism. Results are averaged over 10 runs.
 Implementation & Training Details
@@
 Rest 1 is the only dataset that our model reaches the second-best.
-We attribute this imperfection to the discrepancy between validation and test datasets.
-This is because, through the inspection of the training/validation set performances, we find our model performs better than BrainNetTF on the training and validation dataset.
-More discussions about it can be found in Appendix F.
+We conduct analyses and propose the potential reasons in Appendix F.
+Further, the comparison between our model and the pairwise baseline demonstrates the superiority of incorporating high-order relationships over relying solely on pairwise ones.
 Runtime
-HyBRiD outperforms all other deep learning baselines in efficiency, with $87\%$ faster than the second-fastest one (BrainNetTF). Refer to Appendix I for more runtime details.
+HyBRiD outperforms all other deep learning baselines in efficiency, with $87\%$ faster than the second-fastest one (BrainNetTF). Refer to Appendix J for more runtime details.
 Ablation Studies
-[…]
-Extra Experiments on Synthetic Dataset
-Since there are no ground-truth hyperedges in real-world datasets of learning informative hyperedges towards a specific outcome, we construct a synthetic dataset to verify if our model can recover the correct hyperedge structure under the MIMR objective.
-We use the precision, recall, and F1 score to measure the correctness of the learned hyperedges with respect to the ground truth.
-Although it is challenging to learn hyperedges when only supervised by the task label, we find that our model reaches high performances, with an average improvement of 28.3% in terms of the F1 score, compared to the strongest baselines.
-Details about the synthetic experiments can be found in Appendix G.
+We conduct an ablation study on the effect of our masking mechanism.
+Specifically, we compare our model with $3$ variants: 1) HyBRiDRndMask: Replace the learnable masks with random masks with the same sparsity, initialized at the beginning of training.
+2) HyBRiDNoMask: Do not mask at all, which means all nodes and their features are visible to each head.
+3) HyBRiDSoftMask: Remove the indicator function and use $p^{k}_{\theta,i}$ directly in Equation 3.
+Ablation results are shown in Table 3.
+We find the original HyBRiD and the HyBRiDSoftMask outperform all other variants, which demonstrates the effect of learnable masks.
+Moreover, the original HyBRiD is better than its soft version HyBRiDSoftMask, which demonstrates our sparse and succinct representations preserve better information than smooth ones.
+Other ablation studies such as the choices of the number of hyperedges and choices of $\beta$ can be found in Appendix I.2.
 Figure 4: Hyperedge profiles.
@@
 And the total number of pairwise edges is $26,896$.
+Additional Experiments on Synthetic Dataset
+Since there are no ground-truth hyperedges in real-world datasets of learning informative hyperedges towards a specific outcome, we construct a synthetic dataset to verify if our model can recover the correct hyperedge structure under the MIMR objective.
+We use the precision, recall, and F1 score to measure the correctness of the learned hyperedges with respect to the ground truth.
+Although it is challenging to learn hyperedges when only supervised by the task label, we find that our model reaches high performances, with an average improvement of 28.3% in terms of the F1 score, compared to the strongest baselines.
+Details about the synthetic experiments can be found in Appendix G.
+Additional Experiments on Model Fit
+We further discuss the model’s goodness of fit in Appendix H, with Mean Squared Error (MSE) as the evaluation metric.
+Our model outperforms the state-of-the-art model in 9 out of 11 datasets, with an average improvement of $11.9\%$.
 5.2 Further Analysis
@@
 Hyperedges with Higher Degree are More Significant
-Since CPM conducts a significance test on pairwise edges and hyperedges internally based on a linear regression model, we can obtain a P-value for each hyperedge from the significance test.
+Since CPM conducts a significance test (details can be found in Appendix B) on pairwise edges and hyperedges internally based on a linear regression model, we can obtain a P-value for each hyperedge from the significance test.
 We define the significance of a hyperedge as $1-P_{v}\in[0,1]$ where $P_{v}$ is the P-value of that hyperedge.
-The relationship between hyperedge degree and its significance is shown in Figure  3(b).
+The relationship between hyperedge degree and its significance is shown in Figure  4(b).
 In this figure, we find a strong positive correlation between a hyperedge’s degree and its significance, which indicates that interactions of multiple brain regions play more important roles in cognition than pairwise or individual ones.
 It is also worth mentioning that there is a turning point around degree $5$, which corresponds to the valley around $5$ in Figure 4(a).
-Comparison with Pairwise Edges
+High-order relationships are Better than Pairwise Ones
 To compare the significance in cognition between pairwise edges and learned hyperedges, we plot the number of remaining edges under different thresholds in Figure 4(c).
@@
 This evidence shows that our high-order relationships are much more significant than the traditional pairwise connectivity, which implies relationships involving multiple brain regions could be much more essential in cognition.
-Region Importance
-Figure 5: Visualization of the frequency of each region under different fMRI tasks.
-To better understand the roles of each brain region in cognition under different fMRI tasks, we study the frequency at which each region appears in a hyperedge out of all identified hyperedges.
-The frequency, which can be considered as a measure of region importance, is visualized in Figure 5.
-Visual regions (Fusiform, Cuneus, Calcarine) are especially active due to the intensive visual demands in all three tasks.
-We found that the Inferior frontal gyrus, opercular part and the Anterior cingulate cortex, pregenual, recognized for their participation in response inhibition (Pornpattananangkul et al., 2016), frequently appear in the SST task.
-This aligns with what the SST task was designed to test.
-Interestingly, of the three tasks (SST, EN-back, MID), only EN-back prominently involves the Amygdala, a key region for emotion processing.
-This makes sense as EN-back is the only task related to emotion.
-More visualizations of individual hyperedges can be found in Appendix J.
+Hyperedge Case Study
+We visualize the most significant hyperedge of the EN-back task in Figure 5.
+We observe a coordinated interaction of numerous brain regions, each fulfilling specific roles.
+Notably, some of these regions serve multi-functional purposes:
+Figure 5: Visualization of the most significant hyperedge of the EN-back task.
+  • Memory Processing ParaHippocampal_L, Temporal Mid: Essential for memory encoding and retrieval, these regions are integral to the EN-back task, facilitating the recall of previously viewed images.
+  • Emotional Processing Amygdala_R: The amygdala is crucial for the processing of emotions, such as fear and pleasure.
+Since the EN-back task involves emotional stimuli, it is reasonable that the region is connected by the hyperedge.
+  • Visual Processing: Calcarine_R, Lingual_L, Fusiform_L.
+These regions are responsible for visual perception and some of them are related to complex visual contents like symbols and human faces, which were presented during the fMRI task.
+  • Sensory SupraMarginal_L: It is responsible for interpreting tactile sensors and perceiving limbs location.
+Its involvement is likely due to the requirement for participants to engage in specific physical actions, such as pressing buttons, during the task.
+Temporal Mid: It functions in multi-modal sensory integration.
+  • Motor Control Cerebellum: It is primarily responsible for muscle control.
+Caudate: It plays a crucial role in motor processes.
+Its involvement is likely attributed to participants engaging in physical actions, like pressing buttons.
+  • Cognitive Control ACC_pre_L: In the EN-back task, this region is likely crucial for maintaining focus, error detection and correction, conflict management in working memory, and modulating emotional responses to the task’s demands.
+More visualizations about individual region importance can be found in Appendix K.
 6 Conclusion
@@
 Given that fMRI tasks are dynamic, including temporal changes and interactions, it will be interesting to study the evolution of these high-order relationships.
-7 Impact Statements
+Additionally, HyBRiD does not offer a method for interpreting complex high-order relationships.
+This limitation is not specific to HyBRiD, but is a common challenge in analyzing such relationships.We propose a hierarchical strategy that has the potential to interpret them to a certain extent, which is detailed in Appendix L
+Impact Statement
 This paper presents work whose goal is to advance the field of Machine Learning.
@@
 The statistics of the number of instances and the time series length are summarized in Table 4.
-Dataset ABIDE ABCD
-FIQ VIQ PIQ Rest 1 SST 1 EN-back 1 MID 1 Rest 2 SST 2 EN-back 2 MID 2
-#(instances) 1035 1035 1035 1676 1673 1678 1678 1949 1053 1044 1062
-length of time series 196 196 196 375 437 362 403 375 437 362 403
 Table 4: Statistics of datasets we use.
@@
 For example, Rest 1 means resting-state fMRI data from timepoint $1$.
+Dataset ABIDE ABCD
+FIQ VIQ PIQ Rest 1 SST 1 EN-back 1 MID 1 Rest 2 SST 2 EN-back 2 MID 2
+#(instances) 1035 1035 1035 1676 1673 1678 1678 1949 1053 1044 1062
+length of time series 196 196 196 375 437 362 403 375 437 362 403
 Appendix E Training Details
@@
 See Table 5 for the software we used and the versions.
+Table 5: Software versions.
 software version
@@
 nibabel 4.0.2
-Table 5: Software versions.
 Hyperparameter Choices
-The hyperparameters selection is shown in Table 6. Some crucial hyperparameters ablation experiments can be found in Appendix H.
-notation
-meaning
-value
-$lr$
-learning rate
-$1\times 10^{-3}$
-$K$
-number of hyperedges
-$32$
-$\beta$
-trade-off coefficients information bottleneck
-$0.2$
-$[h_{1},h_{2},h_{3}]$
-hidden sizes of the dim reduction MLP
-$[32,8,1]$
-$B$
-batch size
-$64$
+The hyperparameters selection is shown in Table 6. Some crucial hyperparameters ablation experiments can be found in Appendix I.
 Table 6: Hyperparameter choices.
+notation meaning value
+$lr$ learning rate $1\times 10^{-3}$
+$K$ number of hyperedges $32$
+$\beta$ trade-off coefficients information bottleneck $0.2$
+$[h_{1},h_{2},h_{3}]$ hidden sizes of the dim reduction MLP $[32,8,1]$
+$B$ batch size $64$
 Appendix F Why is HyBRiD not the Best on the Rest 1 Dataset?
-Our model reaches the second-best on the Rest 1 dataset.
-To find out the reason, we inspect the performance on both the training, validation and test datasets.
-We summarize the performances of our model and the SOTA one in the table below
-model train validation test
-BrainNetTF $0.967$ $0.484$ $0.334$
-HyBRiD (Ours) $0.984$ $0.491$ $0.223$
-Table 7: Our model’s performance compared to the SOTA one on the Rest 1 dataset.
-We find our model outperforms BrainNetTF on both training and validation datasets but fails on test datasets.
-Therefore, given that Rest 1 is the most noisy dataset, we argue that the imperfection is because
-1) Our model overfits the training dataset more than BrainNetTF.
-2) The validation score is higher, indicating the reason is the discrepancy between the training dataset and the test dataset.
+At present, we do not fully understand the reasons for the phenomena observed. The hypotheses outlined below are based on our preliminary observations and analyses.
+Hypothesis 1: Sparsity in Resting-State Connectivity
+We propose that resting-state connectivity is sparser and more diffuse compared to task-based connectivity.
+This assertion is refined from our initial claim that the connectivity relevant to the phenotypic outcomes during the resting state is notably sparser than during the task-based state.
+Upon examining traditional pairwise connections, we summarized the number of significant connections most related to the phenotypic outcome in Table 7.
+Our analysis indicates that Rest 1 exhibits the fewest phenotype-related connections, suggesting that these connections are indeed sparse.
+We did not evaluate high-order relationships due to the lack of ground truth data for such connections.
+Further investigation is required to validate this hypothesis.
+SST $1$ EN-back $1$ MID $1$ Rest $1$ SST $2$ EN-back $2$ MID $2$ Rest $2$
+# (connections) $3164$ $3414$ $2989$ $2036$ $3112$ $3014$ $2692$ $3896$
+Table 7: The number of significant pairwise edges selected by CPM in different tasks and timepoints.
+Hypothesis 2: Comparative Quality of Data Across Timepoints
+We hypothesize that the data quality at timepoint 1 is inferior to that at timepoint 2.
+The performance metrics, as shown in the Table 8, are better at timepoint 2 than at timepoint 1.
+Moreover, timepoint 1 exhibits more severe overfitting compared to timepoint 2.
+This discrepancy may be attributed to the higher relevance of the information (pertaining to the outcome) and lower noise levels at timepoint 2.
+This hypothesis also requires further verification through detailed investigations and experiments.
+SST $1$ EN-back $1$ MID $1$ Rest $1$ SST $2$ EN-back $2$ MID $2$ Rest $2$
+train $0.988$ $0.984$ $0.984$ $0.984$ $0.982$ $0.977$ $0.978$ $0.976$
+test $0.361$ $0.348$ $0.386$ $0.223$ $0.738$ $0.714$ $0.816$ $0.730$
+Table 8: Performances on training and test dataset in different tasks and timepoints.
+Conclusion
+Brain activities during resting states are not driven by external tasks, leading to more diffuse and less predictable patterns of activation.
+The low data quality of ABCD timepoint 1 (compared to timepoint 2) even intensifies this issue.
+Therefore, high-order relations might not be a good inductive bias on the Rest 1 dataset since the connections might be much more sparse and involve fewer nodes.
 Appendix G Evaluation on Synthetic Dataset
 Dataset Synthesis
+Table 9: Precision, recall and F1 score on synthetic dataset under different $K$, where $K$ is the number of hyperedges.
+The performance of the Random method increases as $K$ increases because the Hungarian algorithm is likely to provide better matches when there are more candidates.
 Metric $K=1$ $K=5$ $K=10$ $K=30$
@@
 HyBRiD (Ours) $\mathbf{1.000}$ $\mathbf{1.000}$ $\mathbf{1.000}$ $\mathbf{0.595}$ $\mathbf{0.680}$ $\mathbf{0.631}$ $\mathbf{0.504}$ $\mathbf{0.604}$ $\mathbf{0.549}$ $\mathbf{0.428}$ $\mathbf{0.393}$ $\mathbf{0.401}$
-Table 8: Precision, recall and F1 score on synthetic dataset under different $K$, where $K$ is the number of hyperedges.
-The performance of the Random method increases as $K$ increases because the Hungarian algorithm is likely to provide better matches when there are more candidates.
 Our synthetic dataset is constructed as follows:
@@
 This demonstrates that our model can learn the hyperedges well under the MIMR objective, without the direct supervision of hyperedge ground truths.
-Appendix H More Ablation Studies
-H.1 Choices of Number of Hyperedges $K$
+This also inspires us to use automated model evaluation techniques (Wang et al., 2023a; Peng et al., 2023, 2024) on real-world data in the future where labelled high-order relationships are not readily accessible.
+Appendix H Model Fit Performance
+We report the goodness of fit of our model and the state-of-the-art baseline in Table 10, with the Mean Square Error (MSE) as the metric.
+Model FIQ VIQ PIQ
+BrainNetTF $5.917$ $5.429$ $3.355$
+HyBRiD (ours) $\mathbf{3.477}$ $\mathbf{4.331}$ $\mathbf{2.806}$
+Table 10: Performance of the model and baselines (MSE) in fitting the target on the ABIDE dataset.
+Model SST 1 EN-back 1 MID 1 Rest 1 SST 2 EN-back 2 MID 2 Rest 2
+BrainNetTF 1.153 1.373 $\mathbf{1.147}$ $\mathbf{1.244}$ 0.616 0.677 0.720 0.750
+HyBRiD (ours) $\mathbf{1.031}$ $\mathbf{1.348}$ $1.275$ $1.470$ $\mathbf{0.604}$ $\mathbf{0.641}$ $\mathbf{0.543}$ $\mathbf{0.464}$
+Table 11: Performance of the model and baselines (MSE) in fitting the target on the ABCD dataset.
+Appendix I More Ablation Studies
+I.1 Choices of Number of Hyperedges $K$
 As explained in Section 4, we use $K$ heads for $K$ hyperedges. We study the correlation between the $r$ value and the number of hyperedges on three datasets:
@@
 To improve the efficiency while ensuring the performance, we choose $K=32$.
-H.2 Choices of the Trade-off Coefficient $\beta$
+I.2 Choices of the Trade-off Coefficient $\beta$
 In our optimization objective 8, $\beta$ acts as a trade-off parameter, which is a non-negative scalar that determines the weight given to the second term relative to the first.
@@
 Accordingly, we adopt $\beta=0.2$.
-Appendix I Runtime Comparison
+Appendix J Runtime Comparison
 Figure 9 summarizes the per-batch training time of all deep learning models.
@@
 Figure 9: Training time per batch of all deep learning models.
-Figure 10: Visualization of the most significant hyperedge of the EN-back task.
-Appendix J More Visualizations
+Appendix K More Visualizations
+Task-Based Brain Region Importance
+Figure 10: Visualization of the frequency of each region under different fMRI tasks.
+To better understand the roles of each brain region in cognition under different fMRI tasks, we study the frequency at which each region appears in a hyperedge out of all identified hyperedges.
+The frequency, which can be considered as a measure of region importance, is visualized in Figure 10.
+Visual regions (Fusiform, Cuneus, Calcarine) are especially active due to the intensive visual demands in all three tasks.
+We found that the Inferior frontal gyrus, opercular part and the Anterior cingulate cortex, pregenual, recognized for their participation in response inhibition (Pornpattananangkul et al., 2016), frequently appear in the SST task.
+This aligns with what the SST task was designed to test.
+Interestingly, of the three tasks (SST, EN-back, MID), only EN-back prominently involves the Amygdala, a key region for emotion processing.
+This makes sense as EN-back is the only task related to emotion.
 Resting-State Brain Region Importance
@@
 Figure 11: Region importance of resting state.
-Hyperedge Case Study
-We visualize the most significant hyperedge of the EN-back task in Figure 10.
-We observe a coordinated interaction of numerous brain regions, each fulfilling specific roles.
-Notably, some of these regions serve multi-functional purposes:
-  • Memory Processing ParaHippocampal_L, Temporal Mid: Essential for memory encoding and retrieval, these regions are integral to the EN-back task, facilitating the recall of previously viewed images.
-  • Emotional Processing Amygdala_R: The amygdala is crucial for the processing of emotions, such as fear and pleasure.
-Since the EN-back task involves emotional stimuli, it is reasonable that the region is connected by the hyperedge.
-  • Visual Processing: Calcarine_R, Lingual_L, Fusiform_L.
-These regions are responsible for visual perception and some of them are related to complex visual contents like symbols and human faces, which were presented during the fMRI task.
-  • Sensory SupraMarginal_L: It is responsible for interpreting tactile sensors and perceiving limbs location.
-Its involvement is likely due to the requirement for participants to engage in specific physical actions, such as pressing buttons, during the task.
-Temporal Mid: It functions in multi-modal sensory integration.
-  • Motor Control Cerebellum: It is primarily responsible for muscle control.
-Caudate: It plays a crucial role in motor processes.
-Its involvement is likely attributed to participants engaging in physical actions, like pressing buttons.
-  • Cognitive Control ACC_pre_L: In the EN-back task, this region is likely crucial for maintaining focus, error detection and correction, conflict management in working memory, and modulating emotional responses to the task’s demands.
+Appendix L Discussion of the Possible Methods to Interpret Hyperedges
+As mentioned in the main text, high-order relationships are much harder to interpret than pairwise ones given the exponential complexity.
+Here we propose a potential hierarchical strategy that tries to interpret them.
+  • Adaptive Clustering Algorithm with Adjustable Granularity: Suppose we have a clustering algorithm, where the granularity (or number of clusters) can be controlled.
+The clustering algorithm can cluster brain regions based on their functions (e.g.
+motor, visual, …) to different function modules (Shen et al., 2010).
+  • Initial Analysis at Lower Granularity: Set the granularity at a low level (i.e.
+small number of clusters).
+In this case, the high-order relationships are much easier to interpret (since the degree is low).
+However, some high-order relationships will connect to the same set of clusters, thus implying the interactions between the same set of function modules.
+For example, high-order relations $h_{1}$ and $h_{2}$ both connect to visual, motor, and emotion.
+  • Refinement through Increased Granularity: Find the high-order relationships that connect to the same set of function modules, and increase the granularity level, so we can tell the difference between them.
+For example, $h1$ may connect to the left part of the visual module, while $h_{2}$ connects to the right part.